### v1

# Keeping Up with the Models: Online Deployment and Routing of LLMs at Scale

Abstract
The rapid pace at which new large language models (LLMs) appear—and older ones become obsolete—forces LLM service providers to juggle a streaming inventory of models while respecting tight deployment capacity and per-query cost budgets. We cast the reality as an online decision problem that couples stage-wise deployment, made at fixed maintenance windows, with per-query routing among the models kept live. We introduce StageRoute, a hierarchical algorithm that (i) optimistically selects up to $M_{\max}$ models for the next stage using reward upper-confidence and cost lower-confidence bounds, then (ii) solves a budget-constrained bandit sub-problem to route each incoming query. We prove that StageRoute achieves a regret of $\tilde{\mathcal{O}}(T^{2/3})$ and provide a matching lower bound, thereby establishing its near-optimality. Moreover, our experiments confirm the theory, demonstrating that StageRoute performs close to the optimum in practical settings.
1 Introduction
The proliferation of large language models (LLMs) has revolutionized a broad array of applications, delivering unprecedented advances in natural-language understanding and generation [35, 11, 44, 1, 14, 40]. Yet LLMs differ greatly in both performance and cost: some offer state-of-the-art capabilities at a premium, while others are more affordable but less effective. Practitioners must therefore strike a careful balance between accuracy and expenditure when deploying models. This need has given rise to the problem of LLM routing [18, 24], in which a system decides—query by query—which model to invoke, aiming to maximize task quality subject to real-time cost constraints.
The LLM ecosystem is evolving at breakneck speed: new models appear continuously, each with its own accuracy–latency profile and cost structure (for example, per-token API pricing) [21]. The challenge is even sharper in production settings, where the pool of available models changes over time and operators must honor additional constraints such as throughput limits and deployment quotas. Azure OpenAI Service, for instance, caps each resource at 32 standard deployments and 5 fine-tuned deployments by default, and enforces model-specific rate ceilings—GPT-4.1 is limited to 1,000 requests per minute (RPM) and 1 million tokens per minute (TPM) [32].
Motivated by these challenges, we introduce an online decision framework that tackles two intertwined tasks (see Figure 1 for an overview): (1) Stage-wise deployment: At fixed update intervals, decide which LLMs to add or retire, accounting for evolving prices and non-trivial initialization or licensing fees; and (2) Per-query routing: Between updates, assign each user query to a currently deployed model to maximize quality while honoring strict cost budgets. Unlike approaches that assume a static model pool or rely on fully offline retraining, our framework explicitly allows new LLMs to arrive over time. Within each update window it can (i) invest the one-time cost to deploy a promising newcomer discovered during exploration, or (ii) drop a persistently underperforming or overpriced model. This adaptability lets service providers keep pace with rapid LLM advances, maintain service-level quality, and stay within tight resource budgets.
This setting is challenging on several fronts. First, because new—and potentially superior—models continuously appear, any algorithm must adapt its strategy as the candidate pool evolves. Second, updates to the deployed model set occur only at fixed intervals, creating a hierarchical decision structure: long-horizon deployment choices determine the tactical, per-query routing possibilities between updates. Finally, the problem combines a global cost budget, per-model capacity limits, and the combinatorial task of selecting a small subset of models under uncertainty—dramatically increasing its complexity. Standard online-learning tools such as multi-armed bandits or online convex optimization do not natively handle this mix of dynamic availability, staged deployment, and coupled budget-and-capacity constraints.
To tackle these challenges, we introduce StageRoute, an algorithm that mirrors the problem’s hierarchical structure (see Figure 1). At each update boundary, StageRoute enters a model deployment phase, selecting the set of models for the next stage. Using optimistic upper-confidence bounds (UCB) for rewards and conservative lower-confidence bounds (LCB) for costs, it picks the subset that maximizes expected utility while respecting the global budget, per-model capacity limits, and the $M_{\max}$ deployment cap. During the ensuing stage, every incoming query triggers StageRoute’s request routing phase: a linear program over the currently deployed models produces a routing distribution that maximizes estimated quality under the same cost and capacity constraints, and the query is dispatched by sampling from this distribution. This two-level design links strategic stage-level deployment to fine-grained, adaptive routing, allowing the system to integrate new information both across and within stages.
In summary, our contributions in this paper are threefold:
$\bullet$ Problem formulation and algorithm. We formalize the online LLM routing problem with dynamic arrivals and staged deployment and present StageRoute, a hierarchical algorithm that decouples strategic deployment from per-query routing via confidence bounds.
$\bullet$ Theoretical guarantees. We upper bound StageRoute’s regret against an offline oracle and prove a matching $\Omega(T^{2/3})$ lower bound for any online algorithm.
$\bullet$ Empirical evaluation. Experiments show that StageRoute matches a strong oracle within tight budgets and stays robust across key parameters settings.
Figure 1: Overview of the StageRoute workflow. Newly released LLMs (green) continuously enter the candidate pool. At each scheduled update point, StageRoute selects up to $M_{\max}$ to deploy (blue). Between updates, every incoming query is routed among the currently deployed models (orange). This two-level loop lets StageRoute assimilate fresh models, stay within cost constraints, and adapt routing decisions in real time.
2 Related Work
LLM Routing. The central aim of LLM routing is to strike the best balance between task performance (e.g., response quality or accuracy) and operational metrics such as cost and latency [18, 4]. Existing work follows three main architectural patterns. Ensemble strategies query multiple models in parallel to boost robustness, at the expense of higher cost and latency [42, 26]. Cascade strategies issue queries sequentially—typically starting with a cheaper model and escalating only when necessary—thereby reducing cost but sometimes increasing latency [12, 22, 45, 4].Direct-routing strategies train a policy or classifier that selects a single LLM per query [34, 21, 47, 50]. Benchmark suites such as RouterBench [24] facilitate systematic comparison of these approaches. Related work on mixture-of-experts (MoE) models explores routing within a single large model [19, 20, 36]. More recently, bandit formulations have been applied to static LLM routing [43, 17, 31, 33]. Most prior studies, however, assume a fixed set of available models and focus solely on per-query decisions. In contrast, we explicitly model dynamic LLM arrivals, enforce staged deployment updates—where the active model set can change only at periodic intervals—and respect tight constraints on budget, throughput, and the maximum number of simultaneously deployed models ($M_{\max}$).
Multi-armed Bandits. Our formulation builds on the multi-armed-bandit (MAB) paradigm, in which an agent sequentially selects among several actions (“arms”) with initially unknown rewards, learning to maximize its cumulative payoff through exploration and exploitation [7, 39]. Two MAB extensions are especially pertinent: (i) Bandits with knapsacks (BwK). Here each arm pull yields a reward and consumes limited resources from one or more budgets; the objective is to maximize total reward without overspending [10, 5, 25, 28]. Our long-term cost constraint fits naturally into this framework. (2) Streaming bandits. In this setting new arms arrive over time—often under memory or attention limits—so the agent must adapt to a continually expanding action set [6, 27, 3, 41, 30, 38, 49]. The steady appearance of new LLMs places our problem squarely in this category. Our scenario combines both challenges and introduces an additional wrinkle: staged deployment. The set of usable LLMs (arms) can change only at discrete update points, creating a hierarchical decision process in which infrequent, strategic deployment choices constrain frequent, tactical routing decisions. This structure is largely absent from existing BwK or streaming-bandit literature and motivates the two-level algorithm we develop.
3 System Model
We study an online LLM routing framework: at each round $t$, a query arrives and must be routed to a suitable large language model (LLM). New models may appear at any time, yet they can be activated only at discrete deployment intervals (Figure 1). We describe each component of the model in detail.
Rounds (Queries). Let $T$ denote the total number of user queries, indexed by $[T]=\{1,2,\dots,T\}$. At each time step $t\in[T]$, a query arrives and is immediately rerouted to an LLM $m_{t}$ chosen from the currently deployed models according to the algorithm’s policy.
LLM Pool and Deployment Schedule. Let $\mathcal{M}_{t}$ be the set of all LLM models that exist and could, in principle, be deployed by time $t$. Each model $m$ has an availability time $t_{m}$; hence $m\in\mathcal{M}_{t}$ exactly when $t\geq t_{m}.$ Deployment changes occur only at fixed intervals. Divide the horizon of $T$ queries into $K$ equal-length stages of $T/K$ rounds (assume $T$ is divisible by $K$). The start of stage $k$ is $\tau_{k}=(k-1)T/K+1,k=1,\dots,K.$ Let $M_{\max}$ be the maximum number of models that the algorithm can deploy simultaneously. At each update point $\tau_{k}$, the algorithm $\mathcal{A}$ selects a deployed set $\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$ and $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$, which then remains fixed for all $t\in[\tau_{k},\tau_{k+1}).$ Queries arriving during that stage must be routed to models in the active set $\mathcal{D}_{k}(\mathcal{A})$. Thus $\mathcal{M}_{t}$ is the available pool at time $t$, while $\mathcal{D}_{k}(\mathcal{A})$ is the active subset that can actually serve queries during stage $k$.
Operational Performance and Cost. Each model $m$ has an unknown performance distribution $\nu_{m}(\cdot)$ supported on $[0,1]$. When $m$ is selected at time $t$, the observed score $r_{t}\in[0,1]$ is drawn from $\nu_{m}(\cdot)$ with mean $\mu_{m}\;=\;\mathbb{E}_{x\sim\nu_{m}(\cdot)}[x].$ Invoking model $m$ on a query also incurs an operational cost $c_{m_{t}}$ that combines:
(i) Input Cost: $c_{m_{t}}^{(\mathrm{in})}=\bigl{(}\text{\# tokens of input at time }t\bigr{)}\times p_{\mathrm{in}},$ where $p_{\mathrm{in}}$ is the per-token input price;
(ii) Output Cost: $c_{m_{t}}^{(\mathrm{out})}=\bigl{(}\text{\#tokens of response by }m\text{ for query }t\bigr{)}\times p_{\mathrm{out}},$ with output length drawn from a model-specific distribution $\xi_{m}(\cdot)$ and unit price $p_{\mathrm{out}}$.
Thus the total cost is $c_{m_{t}}=c_{m_{t}}^{(\mathrm{in})}+c_{m_{t}}^{(\mathrm{out})}.$ Because the output token count—and therefore $c_{m_{t}}^{(\mathrm{out})}$—depends on the specific model and query and is sampled from $\xi_{m}(\cdot)$, the total operational cost $c_{m_{t}}$ for a query handled by model $m_{t}$ is itself a random variable. Operational cost $c_{m_{t}}$ is inherently bounded by per-token pricing and practical limits on sequence length and generation (e.g., context-window and token caps). Hence we assume $c_{m_{t}}\in[c_{1},c_{2}]$ for known constants $0<c_{1}\leq c_{2}<\infty$.
Cost Constraints and Capacity Limits. When a query arrival at time $t$, with stage index $k$ such that $\tau_{k}\leq t<\tau_{k+1}$, the algorithm $\mathcal{A}$ selects a model $m_{t}\in\mathcal{D}_{k}(\mathcal{A})$. It then observes the reward $r_{t}\sim\nu_{m_{t}}(\cdot)$ and incurs cost $c_{m_{t}}$. The goal is to maximize average reward subject to (1) Long-term budget: $\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\bigl{[}\sum_{t=1}^{T}c_{m_{t}}\bigr{]}\leq b.$ (2) Per-model capacity: For each deployed LLM model $m$, we specify a load cap $\alpha_{m}$. Let $p_{t}(m)$ be the probability that the routing policy assigns the query at time $t$ to model $m$. While $t$ lies in stage $k$, $p_{t}(m)\leq\alpha_{m},\forall m\in\mathcal{D}_{k}(\mathcal{A}).$
Here, $p_{t}(m)$ is the probability that the routing policy assigns the query at time $t$ to model $m$. This constraint caps the instantaneous load share each model may receive. For a deterministic decision $m_{t}$, $p_{t}(m_{t})=1$ and $p_{t}(m)=0$ for $m\neq m_{t}$; the requirement then reduces to $\alpha_{m_{t}}\geq 1$. Such limits reflect real-world restrictions—bandwidth, licensing, or throughput—preventing any single model from handling an excessive fraction of queries within a stage. Hence the deployed set’s aggregate capacity must be large enough to serve every arrival, a condition we formalize next.
Assumption 1 (Feasibility).
The constraints are indeed feasible. At every time $t$, one can select a subset $\mathcal{S}\subseteq\mathcal{M}_{t}$ with $|\mathcal{S}|\leq M_{\max}$ such that $\sum_{m\in\mathcal{S}}\alpha_{m}\geq 1$. The budget $b$ is large enough to admit non-trivial routing. When required, we assume Slater’s condition holds, guaranteeing strong duality for the associated optimization problems.
Performance Maximization and Regret. The goal is to maximize the expected cumulative performance, $\mathbb{E}[\sum_{t=1}^{T}\mu_{m_{t}}]$ subject to:
$\triangleright$ Deployment Choice: At each update $\tau_{k}$, select a deployed set $\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$ with $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$.
$\triangleright$ Model Selection: For $t\in[\tau_{k},\tau_{k+1})$, choose $m_{t}\in\mathcal{D}_{k}(\mathcal{A})$ using probabilities $p_{t}(m)$ that sum to $1$.
$\triangleright$ Capacity Constraint: Ensure $p_{t}(m)\leq\alpha_{m}$ for every deployed model $m.$
$\triangleright$ Budget Constraint: Maintain $\lim_{T\to\infty}\frac{1}{T}\mathbb{E}[\sum_{t=1}^{T}c_{m_{t}}]\leq b$.
We measure the online policy’s performance against an optimal offline benchmark. The foundation of this benchmark is the Optimal Performance Rate Function, $V(b,\mathcal{S})$. Given any candidate model set $\mathcal{S}$ and per-query budget $b$, let $V(b,\mathcal{S})$ denote the maximum expected reward per query, which serves as an upper bound for any algorithm operating under these constraints:
[…]
$\qquad\qquad 0\leq p(m)\leq\alpha_{m}\text{ for }m\in\mathcal{S},|\text{supp}(p)|\leq M_{\max}\Big{\}}.$
Here, $\Delta(\mathcal{S})$ is the set of probability distributions over $\mathcal{S}$; $p(m)$ is the probability of selecting model $m$; and $\mathbb{E}[c_{m}]$ is its expected cost. The support constraint $|\text{supp}(p)|\leq M_{\max}$ limits the number of models with positive probability to $M_{\max}$. If no feasible distribution exists—or if $(\mathcal{S}=\varnothing)$—we set $V(b,\mathcal{S})=0$.
The time-varying offline optimum is $\text{OPT}^{*}=\sum_{k=1}^{K}(\tau_{k+1}-\tau_{k})\cdot V(b,\mathcal{M}_{\tau_{k}}),$ where $(\tau_{k+1}-\tau_{k})$ is the length (number of queries) of stage $k$. The regret of an online policy $\mathcal{A}$ is
$\text{Regret}(\mathcal{A})=\text{OPT}^{*}-\mathbb{E}\biggl{[}\sum_{t=1}^{T}\mu_{m_{t}}\biggr{]},$ (2)
i.e., the expected performance gap between $\mathcal{A}$ and the clairvoyant benchmark, where the expectation is over the algorithm’s random choices and outcome variability.
4 StageRoute: Stage-Based LLM Deployment and Routing Algorithm
We now introduce StageRoute (Algorithm 1), a two-level hierarchy that unifies deployment and per-query routing in one algorithm: (i) Strategic layer. At each discrete update point $\tau_{k}$, the algorithm decides which models to deploy, adapting to newly available LLMs while respecting capacity and budget limits. (ii) Tactical layer. Between updates, it routes every incoming query in real time among the currently deployed models. The system starts with prior parameter estimates, an empty unexplored-model list, $\tau_{0}=0$, and an empty initial deployment $\mathcal{D}_{0}(\mathcal{A})$. It then proceeds through $K$ stages defined by update points $\{\tau_{1},\dots,\tau_{K}\}$. At the start of each state $k$ ($t=\tau_{k}$), the algorithm executes two phases in sequence: model deployment followed by request routing.
Model Deployment Phase (Stage Start): At each update point $\tau_{k}$, StageRoute first adds every LLM that has become eligible since the previous update—those with $t_{m}\leq\tau_{k}$ and $t_{m}>\tau_{k-1}$—and initializes their parameter estimates. With the enlarged pool $\mathcal{M}_{\tau_{k}}$, StageRoute then solves the deployment optimization in Eq. (3) to pick the models for stage $k$.
[…]
$0\leq d_{m}\leq\alpha_{m}\text{ for }m\in\mathcal{M}_{\tau_{k}},|\text{supp}(d)|=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|)\Big{\}}.$
The decision vector $d^{*}=(d^{*}_{m})_{m\in\mathcal{M}_{\tau_{k}}}$ maximizes an optimistic performance surrogate while respecting budget and capacity constraints, using upper confidence bounds (UCBs) on rewards ($\mu^{U}_{m}$) and lower confidence bounds (LCBs) on costs ($c^{L}_{m}$), both derived from data up to $\tau_{k}.$
Specifically, let $\bar{\mu}_{m}(\tau_{k})$ and $\bar{c}_{m}(\tau_{k})$ be the empirical mean reward and cost of model $m$ based on $N_{m}(\tau_{k})$ selections observed up to $\tau_{k}$. Define the UCBs and LCBs as:
$\mu^{U}_{m}$ $:=\text{proj}_{[0,1]}\left(\bar{\mu}_{m}(\tau_{k})+2f_{rad}(\bar{\mu}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)\right),$ (4)
$c^{L}_{m}$ $:=\text{proj}_{[c_{1},c_{2}]}\left(\bar{c}_{m}(\tau_{k})-2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)\right).$ (5)
where $\text{proj}_{[a,b]}$ is a projection function onto the interval $[a,b]$ and $f_{rad}(v,n)=\sqrt{\frac{\gamma v}{n}}+\frac{\gamma}{n}$ (for some $\gamma>0$) is a confidence radius function.
The deployment optimization in Eq. (3) maximizes expected utility subject to the budget $b$, per-model capacity limits $\alpha_{m}$, and the cardinality cap $|\text{supp}(d)|=\min(M_{\max},|\mathcal{M}_{\tau_{k}}|)$. Its solution $d^{*}$ determines the active set for stage $k$:
$\mathcal{D}_{k}(\mathcal{A})\leftarrow\{m\in\mathcal{M}_{\tau_{k}}\mid d^{*}_{m}>0\}.$ (6)
The value $d^{*}_{m}$ themselves are not used as routing probabilities; they serve only to pick the most promising feasible models. The deployment set $\mathcal{D}_{k}(\mathcal{A})$ then remains fixed until the next update point $\tau_{k+1}.$
Algorithm 1 StageRoute: Stage-Based LLM Routing
0:  Update points (stage boundaries) $\{\tau_{1},\dots,\tau_{K}\}$; overall operational budget $b$; maximum number of concurrently deployed models $M_{\max}$
1:  Initialize: Prior estimates for their parameters; unexplored set $\mathcal{U}\leftarrow\emptyset$; initial time $\tau_{0}\leftarrow 0$; initial deployed set $\mathcal{D}_{0}(\mathcal{A})\leftarrow\emptyset$
2:  ///// /Stage-wise Deployment Phase:
3:  for each stage $k=1$ to $K$ do
4:     Add newly available models $\{m\mid t_{m}\leq\tau_{k}\text{ and }t_{m}>\tau_{k-1}\}$, and initialize their parameters
5:     Determine the deployment decision vector $d^{*}=(d^{*}_{m})_{m\in\mathcal{M}_{\tau_{k}}}$ by solving DeployOPT (3)
6:     Update the set of deployed models for stage $k$: $\mathcal{D}_{k}(\mathcal{A})\leftarrow\{m\in\mathcal{M}_{\tau_{k}}\mid d^{*}_{m}>0\}$
7:     ///// /Per-query Routing Phase (for query at time $t$):
8:     for each time step $t$ from $\tau_{k}$ to $\tau_{k+1}-1$ (if $k<K$), or to end of horizon (if $k=K$do
9:        Compute the optimal routing distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ by solving RouteOPT (7)
10:        Sample model $m_{t}\sim p_{t}^{*}$, and route request to the chosen model $m_{t}$
11:        Observe the actual reward $r_{t}$ and actual cost $c_{t}$ resulting from using model $m_{t}$
12:        Update the parameters for model $m_{t}$
13:     end for
14:  end for
Request Routing Phase (Intra-Stage): After each deployment update, StageRoute handles each query at time $t\in[\tau_{k},\tau_{k+1}-1)$ (or until $T$ if $k=K$) in four steps as follows: (1) Routing LP. It solves the linear program (LP) in Eq. (7) to compute the optimal routing distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ over the currently deployed models $\mathcal{D}_{k}(\mathcal{A})$.
[…]
$\quad\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}p_{t}(m)=1,0\leq p_{t}(m)\leq\alpha_{m}\text{ for }m\in\mathcal{D}_{k}(\mathcal{A})\Big{\}}.$
This LP maximizes the expected reward by combining the current UCBs for reward ($\mu^{U}_{m}$) and LCBs for cost ($c^{L}_{m}$) while enforcing the per-query budget $b$ (through the cost bounds) and each model’s capacity limit $\alpha_{m};$ the cardinality cap $M_{\max}$ is already respected because $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$. (2) Model selection. A model $m_{t}$ is sampled according to $p_{t}^{*}$ and used to serve the query. (3) Feedback. StageRoute observes the realized reward $r_{t}$ and cost $c_{t}$. (4) Update statistics. StageRoute then updates $\mu_{m_{t}}$, $c_{m_{t}}$ and $N_{m_{t}}$, recomputing the corresponding confidence bounds so that future routing—and the next deployment decision—reflect the latest information.
Following the deployment decision, the algorithm enters an inner loop that processes each query arriving at time $t$ from $\tau_{k}$ to $\tau_{k+1}-1$ (or to the end of the horizon if $k=K$). For each query, the algorithm performs the following steps: It determines the optimal routing probability distribution $p_{t}^{*}=(p_{t}^{*}(m))_{m\in\mathcal{D}_{k}(\mathcal{A})}$ over the currently deployed models $\mathcal{D}_{k}(\mathcal{A})$ by solving the linear program (LP) in Equation (7). This LP maximizes the expected reward based on current UCBs for reward ($\mu^{U}_{m}$) and LCBs for cost ($c^{L}_{m}$) for models in $\mathcal{D}_{k}(\mathcal{A})$. These UCB/LCB estimates are based on all information gathered up to the previous time step $t-1$ and are updated as new observations are made. The routing LP adheres to the long-term budget $b$ (applied on a per-query basis using LCBs for costs) and individual model capacity constraints $\alpha_{m}$ for $m\in\mathcal{D}_{k}(\mathcal{A})$. Note that the $M_{\max}$ constraint is not part of this LP, as the set $\mathcal{D}_{k}(\mathcal{A})$ already satisfies $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$ (due to the constraint in Equation 3 and the formation of $\mathcal{D}_{k}(\mathcal{A})$). A model $m_{t}$ is then sampled according to this distribution $p_{t}^{*}$ (Line 10), and the incoming query is routed to $m_{t}$. The system observes the actual performance (reward) $r_{t}$ and cost $c_{t}$ from the selected model $m_{t}$. Finally, this feedback is used to update the parameters for model $m_{t}$, such as its empirical average reward, empirical average cost, and usage count. These updated statistics are then used to refine the $\mu^{U}_{m_{t}}$ and $c^{L}_{m_{t}}$ estimates for subsequent routing decisions and for the deployment decision at the next stage.
5 Theoretical Analysis
We analyze StageRoute (Algorithm 1) by deriving an upper bound on its cumulative regret and a matching lower bound that applies to any online algorithm. Taken together, these results show that StageRoute is near-optimal in the worst case.
5.1 Upper Bound
Theorem 1.
Consider StageRoute running for $T$ queries divided into $K$ stages, with $N=|\mathcal{M}_{T}|$ models that may arrive over time. Set the confidence parameter to $\gamma=\Theta(\log(NT/\delta))$ to obtain overall confidence $1-\delta$. Then StageRoute’s expected regret satisfies
$\textnormal{\text{Regret}}(\textnormal{{StageRoute}})\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+\frac{NT}{M_{\max}K}\right).$
When $K=\Omega(T^{1/3})$ and $M_{\max}=\Omega(N^{2/3})$, we have $\textnormal{\text{Regret}}(\textnormal{{StageRoute}})\leq\tilde{\mathcal{O}}\left(N^{1/3}T^{2/3}\right).$
The theorem’s primary regret bound demonstrates that overall system performance is governed by a trade-off between two key components: The regret bound reveals a tradeoff between two components. The statistical term $\mathcal{O}(\sqrt{M_{\max}KT\log(NT/\delta)})$ reflects the difficulty of learning optimal routing and deployment. The structural term $\mathcal{O}(NT/(M_{\max}K))$ captures the exploration bottleneck created by probing a large universe of $N$ models through only $M_{\max}$ deployment slots over $K$ stages. This tension means that both the number of stages $K$ and the deployment capacity $M_{\max}$ must be tuned carefully. When chosen well, the overall regret matches the lower bound in Theorem  2, yielding a near-optimal order of $\widetilde{\mathcal{O}}(T^{2/3})$. Although prior work suggests that good routing can rely on a small set of models, our analysis shows this is not true in dynamic environments with continual model turnover. In such settings, maintaining a sufficiently large concurrent deployment capacity is crucial for near-optimal performance.
Our analysis unifies stage-wise decision making with a continually evolving model pool, tackling the practical need to deploy newly released LLMs under tight capacity limits. We expose a new model-discovery bottleneck that captures the exploration challenge created by the mismatch between the limited number of concurrent slots ($M_{\max}$) and the discrete schedule of deployment stages. This bottleneck is distinct from the usual learning-related regret incurred by routing and deployment decisions. The theory shows that near-optimal performance requires setting the stage count to $K=\Theta(T^{1/3})$ and provisioning at least $M_{\max}=\Omega(N^{2/3})$ concurrent slots. These prescriptions offer actionable guidance for system designers and can inform future algorithmic and infrastructure choices in real-world LLM-serving environments.
Proof Sketch..
We decompose the algorithm’s total expected regret into two parts, i.e. the expected routing regret, $\mathcal{R}_{\text{routing}}(T)$ and the expected deployment regret, $\mathcal{R}_{\text{deploy}}(T)$, so that
$\text{Regret}(\mathcal{A})=\underbrace{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k})\right]}_{\mathcal{R}_{\text{deploy}}(T)}+\underbrace{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{D}_{k})-r_{t}\right]}_{\mathcal{R}_{\text{routing}}(T)}.$
The routing regret, $\mathcal{R}_{\text{routing}}(T)$, arises because the algorithm relies on estimated model parameters when routing queries within each deployed set $\mathcal{D}_{k}$. For a single stage $k$, the expected routing regret is bounded by $\mathcal{O}\left(\sqrt{{\gamma}|\mathcal{D}_{k}|T\cdot V(b,\mathcal{D}_{k})/K}\right)$. Summing the regret bounds over all $K$ stages, the total expected routing regret is bounded by
$\mathcal{R}_{\text{routing}}(T)\leq\mathcal{O}\left(\sqrt{\gamma M_{\max}KT}\right).$
The deployment regret, $\mathcal{R}_{\text{deploy}}(T)$, is more intricate and has two component. (i) Learning error: uncertainty in the model estimates when selecting the deployment set $\mathcal{D}_{k}$. (ii) Model discovery bottleneck: the $M_{\max}$-slot limit can delay evaluation and adoption of newly arrived, high-quality models. The sum of these two terms is bounded by
$\mathcal{R}_{\text{deploy}}(T)\leq\mathcal{O}\left(\sqrt{\gamma T\cdot\min(N,KM_{\max})}+\frac{NT}{M_{\max}K}\right).$
Combining the routing- and deployment-regret bounds and substituting $\gamma=\Theta(\log(NT/\delta))$ gives the overall regret bound stated in the theorem; the full proof appears in Appendix B. ∎
5.2 Lower Bound
Theorem 2.
Consider an online LLM routing problem with $M\geq 2$ models over a horizon of $T$ queries. For any online routing algorithm $\mathcal{A}$, when $T$ is sufficiently large relative to $M$, its worst-case expected regret is bounded below by:
$\textnormal{\text{Regret}}(\mathcal{A})\;\geq\;\Omega(T^{2/3}).$
The $\Omega(T^{2/3})$ lower bound captures the intrinsic difficulty of continually tracking the best model as capabilities evolve. The proof uses a batch-based construction in which the “strong” model changes stochastically from one batch to the next, mirroring real LLM ecosystems where newer, better models apear regularly. An online algorithm must swiftly detect and exploit each transiently optimal model before it is displaced, and this necessity yields the unavoidable $T^{2/3}$ regret rate. Thus, the bound is not an artifact of the analysis but a fundamental limit for stage-based routing under dynamic model arrivals. The result also shows that system parameters—such as how often deployments occur and how many models can run concurrently—must be tuned carefully to approach this performance frontier.
Proof Sketch..
The proof constructs a class of adversarial problem instances designed to demonstrate that any online algorithm must incur significant regret on at least one of them. Specifically, the time horizon $T$ is partitioned into $N_{B}$ batches, each of equal length $\Delta=T/N_{B}$. In each batch $j\in\{1,\dots,N_{B}\}$, one out of $M$ models—denoted $m_{s_{j^{*}}}$—is designated as the strong model, offering an expected Bernoulli reward of $\mu_{strong}^{(j)}=\frac{1}{2}+j\epsilon$. The remaining $M-1$ models are treated as weak models, each yielding a lower expected reward of $\mu_{weak}^{(j)}=\frac{1}{2}+(j-1)\epsilon$. This construction captures a realistic abstraction of the modern LLM landscape, where new, more capable models continuously emerge while older ones remain accessible but less effective. The key statistical challenge lies in the unknown sequence of strong models, which must be discovered online. The reward gap $\epsilon>0$ plays a central role in the difficulty of the problem. To ensure that all mean rewards remain within the valid range $[0,1]$, the parameter $\epsilon$ is chosen such that $N_{B}\epsilon\leq c$ for constant $20<c<1/$ 2.
For a fixed batch $j$, the regret incurred in this batch is given by $R_{j}(P_{j,s_{j}^{*}})=\epsilon(\Delta-\mathbb{E}_{P_{j,s_{j}^{*}}}[N_{j,s_{j}^{*}}])$, where $N_{j,s_{j}^{*}}$ denotes the number of times the algorithm selects model $m_{s_{j}^{*}}$ during batch $j$. Leveraging information-theoretic arguments, which connect the difficulty of identifying the optimal model to the number of observations and the magnitude of the reward gap, we derive a bound on the total expected number of pulls across the set of strong models. This bound captures the algorithm’s limited ability to reliably discover and exploit the strong model within each batch, particularly when $\Delta$ is small and $\epsilon$ is narrow. We then lower bound the expected regret for batch $j$, under the assumption that the strong model index $s_{j}^{*}$ is chosen uniformly at random from $\{1,\dots,M\}$. This leads to the bound $R_{T}\gtrsim T\epsilon\left(\frac{1}{2}-\epsilon\sqrt{\Delta}\right)$. Optimizing the bound by selecting $\epsilon\asymp T^{-1/3}$ and $\Delta\asymp T^{2/3}$, and substituting these into the expression above, yields the desired lower bound. A detailed proof is provided in Appendix C. ∎
6 Experiments
This section presents the empirical evaluation of our proposed StageRoute algorithm, with a focus on its performance under diverse operational settings. All reported results are averaged over 10 independent simulation runs, each consisting of $T=10^{4}$ user queries. The experiments involve solving mixed-integer programming (MIP) subproblems using the academic version of the Gurobi Optimizer (version 12.0.1). All runs are executed on a machine equipped with a 12th Gen Intel(R) Core(TM) i9-12900HX processor.
Table 1: LLM comparison by release date: performance (Arena Score/Elo [13]) and output cost (per 1M tokens).
Model Performance Cost
GPT-4o 1336 \$10.00
Gemini-1.5-pro-002 1302 \$5.00
Yi-Lightning 1287 \$0.14
o1-mini 1304 \$0.60
Llama 3.3 70B Instruct 1257 \$0.40
DeepSeek-R1 1358 \$0.55
Gemini 2.0 Flash 1380 \$1.50
Claude 3.7 Sonnet 1300 \$15.00
Hunyuan-turbos 1296 \$0.28
Deepseek-v3 1373 \$0.28
Llama-4-Maveric 1417 \$0.82
GPT-4.1 1366 \$8.00
Grok-3-preview 1403 \$15.00
o3 by OpenAI 1413 \$40.00
Gemini-2.5-Pro 1446 \$10.00
6.1 Simulation Results
LLMs Pool. We consider a set of 15 LLMs, summarized in Table 1, ordered by their official release dates. For each model, we report both performance metrics  [13] and associated costs. Note that the naming of AI models can be highly nuanced, and multiple versions may exist under a similar label. For example, for Gemini 2.5 Pro, we used data corresponding to the Gemini-2.5-Pro-Preview-05-06 version. Additionally, performance scores may fluctuate over time due to model updates, shifts in the user voting population, or changes in evaluation benchmarks. Similarly, costs may vary across time or across versions. The data reported here corresponds to the specific versions and conditions used in our experiments.
These 15 models are introduced sequentially, with the first 6 available at time $t=1$ and one additional model becoming available every 1,000 rounds until all are accessible. Each model’s performance is normalized to lie within the $[0,1]$ interval. For each model $m$, we set its budget weight parameter $\alpha_{m}=0.4$, except for Yi-Lightning, for which we set $\alpha_{m}=1$. This is because Yi-Lightning is the cheapest model initially available to the algorithm, ensuring the feasibility of the mixed-integer optimization problem in every round. At each round, if the algorithm selects a model, it receives a noisy performance and cost observation, where the noise is sampled from a Gaussian distribution centered around the true value.
Baselines. We compare our StageRoute against two baselines. The first is an oracle that has full knowledge of all relevant performance and cost statistics in advance and always selects the optimal set of models. This serves as an upper bound on achievable performance. The second baseline is a greedy strategy that, at each update point, selects the $M_{\max}$ models with the highest utility, computed as the ratio between the UCB of performance and the LCB of cost. This can in fact be viewed as a variant of a UCB algorithm, where selection is based on the UCB of the utility metric. For both our algorithm and the baselines, we set $\gamma=0.2$ and $b=2$. The cost constraint is enforced as a hard budget.
(a) Performance.
(b) Regret.
Figure 2: Performance with $M_{\max}=6$ and
Update Interval 250 rounds.
(c) 500 rounds.
(d) 1,000 rounds.
Figure 3: Varying update intervals with $M_{\max}$ fixed at 6.
(a) $M_{\max}=3$.
(b) $M_{\max}=10$.
Figure 4: Varying $M_{\max}$ with update interval fixed at 500 rounds.
Figure 5: Model choice for $M_{\max}=3$ (Update interval at 500 rounds).
Overall Performance and Regret. We study the effect of two key hyperparameters on algorithm performance. The deployment update interval determines how frequently the algorithm can revise its set of actively deployed LLMs. We test intervals of 250, 500, and 1000 rounds. The second is maximum number of concurrently deployed models ($M_{\max}$), which limits the size of the active model pool at any time. We evaluate settings of $M_{\max}\in\{3,6,10\}$. Figures 55 illustrate the performance dynamics of the proposed algorithm under various settings. We begin with Figure 5, which shows the performance when $M_{\max}=6$ and the deployment update interval is 250 rounds. As shown in Figure 2(a), StageRoute attains performance that closely matches that of the optimal oracle. Meanwhile, Figure 2(b) shows that StageRoute incurs substantially lower regret than the greedy strategy, highlighting its superior ability to balance exploration and exploitation under uncertainty.
$\triangleright$ Effect of Deployment Update Interval. Figures 2(b), 2(c), and 2(d) illustrate the impact of different deployment update intervals. A shorter interval enables the algorithm to adapt more quickly to newly available LLMs as well as to evolving performance and cost estimates. In contrast, infrequent updates—such as every 1000 rounds, as shown in Figure 2(d)—slow down this adaptation process. This reduced responsiveness can lead to consistently higher regret across all algorithms.
$\triangleright$ Impact of $M_{\max}$. Figures 2(c), 3(a), and 3(b) show that a smaller value of $M_{\max}$ limits the algorithm’s flexibility in composing a diverse routing mix. In contrast, increasing the capacity to $M_{\max}=10$ (Figure 3(b)) allows both StageRoute and the greedy strategy to achieve strong performance. These results suggest that the advantage of StageRoute is most pronounced when deployment capacity is limited, while it remains competitive even in high-capacity regimes. Moreover, Figure 5 offers a concrete example of model selection behavior under $M_{\max}=3$ and an update interval of 500 rounds. The algorithm dynamically alternates between expensive proprietary models and cheaper open-source alternatives, effectively adapting to the evolving LLM landscape and maintaining performance in a cost-aware manner.
6.2 Real-World Dataset Evaluation
We then present experimental results on a real-world dataset. Our evaluation is based on the Router-Bench dataset [24], a comprehensive benchmark comprising 36,497 queries sampled from eight diverse natural language processing (NLP) datasets. These datasets encompass both Chinese and English languages and cover a broad spectrum of core tasks relevant to LLMs. Specifically, the benchmark includes seven established evaluation datasets, along with an additional dataset targeting retrieval-augmented generation (RAG).
The seven core datasets span the following categories: commonsense reasoning, including HellaSwag [46], WinoGrande [37], and ARC Challenge [15]; knowledge-based language understanding, represented by MMLU [23]; open-domain dialogue, using MT-Bench [48]; mathematical reasoning, evaluated through GSM8K [16]; and code generation, based on MBPP [8]. The additional RAG dataset is designed to reflect realistic retrieval-augmented tasks encountered in production environments. For each query in the dataset, responses from 11 distinct LLMs are provided, together with corresponding performance scores and cost measurements.
Table 2: List of LLMs sorted by release date, along with their per-query average performance score and average cost.
Model Avg Performance Avg Cost
claude-instant-v1 0.5900 \$0.001236
claude-v1 0.6480 \$0.005870
claude-v2 0.5116 \$0.006153
meta/llama-2-70b-chat 0.6059 \$0.001337
WizardLM/WizardLM-13B-V1.2 0.5392 \$0.000142
meta/code-llama-instruct-34b-chat 0.5040 \$0.000550
mistralai/mistral-7b-chat 0.4999 \$0.000139
gpt-3.5-turbo-1106 0.6867 \$0.000709
gpt-4-1106-preview 0.8048 \$0.007943
zero-one-ai/Yi-34B-Chat 0.7153 \$0.000558
mistralai/mixtral-8x7b-chat 0.6504 \$0.000414
Figure 6: Cumulative performance score over time on the real-world dataset.
In our experiments, we simulate a total of $T=36,\!497$ rounds. At each round, a query is sampled uniformly at random from the dataset, including its associated prompt. The algorithm then selects a model to serve the query, and subsequently receives a noisy reward signal (or correctness label) along with the corresponding cost.
Initially, only 5 models are available for deployment. Then, for every additional 5,000 queries, one new model becomes eligible for deployment, following the release-date ordering specified above. For the remaining parameters, we set $\gamma=0.2$ and $b=0.0002$. The cumulative reward achieved by different methods under this setting is illustrated in Figure 6. As shown in the figure, our algorithm consistently outperforms the Greedy strategy by an even larger margin compared to the synthetic dataset experiments. This further demonstrates the effectiveness and robustness of our method in real-world scenarios.
Limitations and Future Work
In this paper, we introduced a stage-based framework for online LLM routing that explicitly handles changing model availability under budget, capacity, and deployment constraints. The resulting algorithm balances exploration and exploitation in this challenging setting, and we supported it with provable regret bounds and strong empirical results. Regarding societal impacts, our framework has the potential to facilitate broader adoption of LLMs by lowering deployment costs and improving energy efficiency. However, increased accessibility and proliferation of LLMs may also give rise to negative societal consequences. Several limitations remain. Our reward and cost estimates rely on UCB/LCB statistics derived from past observations, which may be inadequate in scenarios where a model’s performance or pricing varies significantly with the input context. In fact, if the number of deployable models is smaller than the number of distinct context clusters, any routing policy suffers $\Omega(T)$ worst-case regret. Future work can exploit structure in the context space. For instance, if rewards or costs vary linearly with contextual features, more sophisticated contextual bandit techniques may yield tighter regret guarantees and better practical performance. Exploring such structured settings is an appealing direction for further research.
Appendix A Summary of Notation
Table 3: Table of Notation Used in the System Model and Algorithms
Symbol Description
General Parameters & Indices
$T$; $t\in[T]$ Total queries (horizon $T$); $t$ is query index in $[T]=\{1,\dots,T\}$.
$K$; $k\in[K]$ Total deployment stages ($K$); $k$ is stage index.
$\tau_{k}$ Start time step of stage $k$; $k$-th deployment update point, $\tau_{k}=(k-1)T/K+1$.
$\tau_{0}$ Initial time for the algorithm, typically 00.
$\mathcal{A}$ The online LLM routing algorithm/policy.
LLMs: Availability & Deployment
$m$; $t_{m}$ An individual LLM $m$; and its availability time $t_{m}$.
$\mathcal{M}_{t},\mathcal{M}_{\tau_{k}}$ Set of LLMs available at time $t$, and specifically at start of stage $k$.
$M_{\max}$ Maximum number of LLMs that can be simultaneously deployed.
$d=(d_{m})$ Deployment decision variable vector in DeployOPT over $\mathcal{M}_{\tau_{k}}$.
$d^{*}=(d^{*}_{m})$ Optimal deployment decision vector from DeployOPT (Eq. 3) at $\tau_{k}$.
$\mathcal{D}_{0}(\mathcal{A})$ Initial set of deployed LLMs by algorithm $\mathcal{A}$ (typically empty).
$\mathcal{D}_{k}(\mathcal{A})$ Set of LLMs deployed by $\mathcal{A}$ in stage $k$ (derived from $d^{*}_{m}>0$).
LLM Performance & Operational Costs
$\nu_{m}(\cdot),\mu_{m}$ Performance distribution for LLM $m$ (on $[0,1]$) and its mean $\mu_{m}=\mathbb{E}_{x\sim\nu_{m}(\cdot)}[x]$.
$r_{t}$ Realized performance from model $m_{t}$ for query $t$, $r_{t}\sim\nu_{m_{t}}(\cdot)$.
$\bar{\mu}_{m}(\tau_{k}),\mu^{U}_{m}$ Empirical mean performance (from $N_{m}(\tau_{k})$ obs. up to $\tau_{k}$) and UCB for $\mu_{m}$.
$p_{\mathrm{in}},p_{\mathrm{out}}$ Per-token prices for input and output.
$c_{m_{t}}^{(\mathrm{in})},c_{m_{t}}^{(\mathrm{out})};c_{m_{t}}$ Input cost, output cost; and total cost $c_{m_{t}}=c_{m_{t}}^{(\mathrm{in})}+c_{m_{t}}^{(\mathrm{out})}$ for $m_{t}$ on query $t$.
$\xi_{m}(\cdot)$ True (unknown) distribution of output token length for LLM $m$.
$\mathbb{E}[c_{m}]$ True expected operational cost of LLM $m$.
$\bar{c}_{m}(\tau_{k})$ Empirical mean operational cost of LLM $m$ based on $N_{m}(\tau_{k})$ selections up to $\tau_{k}$.
$c^{L}_{m}$ Lower Confidence Bound (LCB) on the expected operational cost $\mathbb{E}[c_{m}]$.
$c_{1},c_{2}$ Fixed lower and upper bounds for any $c_{m_{t}}$, $0<c_{1}\leq c_{2}<\infty$.
Routing & Constraints
$m_{t}$ LLM selected by the algorithm to handle query $t$.
$p_{t}(m)$ Probability assigned by a routing policy to LLM $m\in\mathcal{D}_{k}(\mathcal{A})$ for query $t$.
$p_{t}^{*}=(p_{t}^{*}(m))$ Optimal routing probabilities from RouteOPT (Eq. 7).
$b$ Long-term average operational cost budget per query.
$\alpha_{m}$ Capacity limit (maximum load share / selection probability constraint) for LLM $m$.
$\Delta(\mathcal{S})$ Set of all probability distributions over a set of LLMs $\mathcal{S}$.
Parameter Estimation & Confidence Bounds
$N_{m}(\tau_{k})$ Number of times LLM $m$ has been selected and observed up to $\tau_{k}$.
$f_{rad}(v,n),\gamma$ Confidence radius function $f_{rad}(v,n)=\sqrt{\gamma v/n}+\gamma/n$ (with parameter $\gamma>0$).
$\text{proj}_{[a,b]}(x)$ Projection of value $x$ onto the interval $[a,b]$.
Offline Benchmark & Regret
$V(b,\mathcal{S})$ Optimal Performance Rate Function: max expected performance from set $\mathcal{S}$.
$\text{supp}(p)$ Support of a probability distribution $p$.
$\text{OPT}^{*}$ Expected cumulative reward of the time-varying offline optimal policy.
$\text{Regret}(\mathcal{A})$ Regret of online algorithm $\mathcal{A}$.
Appendix B Technical Analysis
In this section, we analyze the regret of StageRoute (Algorithm 1). Recall that in our setting, let $N=|\mathcal{M}_{T}|$ denote the total number of models that may arrive over the course of the time horizon. We assume that $N$ is significantly larger than $M_{\max}$, the maximum number of models that can be deployed simultaneously. Moreover, the number of update points $K$ is assumed to be much smaller than the total number of queries $T$. These assumptions reflect practical constraints: it is typically infeasible to deploy all available models—including those released in the future—due to resource limitations, and continuously updating the deployed LLM pool in real time is operationally impractical.
B.1 Concentration Inequality
We employ the following standard concentration inequality and related supporting lemmas.
Lemma 1 ([29, 9]).
Consider a sequence of random variables $x_{1},x_{2},\ldots,x_{n}$. Let $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ be the empirical average and $v=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[x_{i}|x_{1},\ldots,x_{i-1}]$ (if $i=1$, the expectation is unconditional). If the values $x_{i}$ are in $[0,1]$ (e.g., performance $r_{t}$, or cost $c_{t}$ under our assumption), then for each $\gamma>0$,
$\mathbb{P}[|v-\bar{x}|\leq f_{rad}(\bar{x},n)\text{ and }f_{rad}(\bar{x},n)\leq 3f_{rad}(v,n)]\geq 1-\exp(-\Omega(\gamma)),$ (8)
where $f_{rad}(v,n)=\sqrt{\frac{\gamma v}{n}}+\frac{\gamma}{n}$. This result also holds if $x_{1},\ldots,x_{n}$ are independent samples from a distribution with mean $v$ and values in $[0,1]$.
For clarity and to simplify the application of concentration inequalities, we assume throughout this analysis that all operational costs $c_{m_{t}}$ are bounded such that $0<c_{1}\leq c_{m_{t}}\leq c_{2}\leq 1$. This ensures that costs, like rewards (which are in $[0,1]$), fall within a $[0,1]$ range (or a sub-interval thereof). This assumption does not affect the order of the regret bounds, as any scaling factors related to a broader cost range would typically be absorbed into the constants within the $\mathcal{O}(\cdot)$ notation.
Lemma 2 ([9], adapted).
Let $\mathcal{D}_{k}(\mathcal{A})$ be the set of deployed models in stage $k$. For any vectors $\boldsymbol{a}=(a_{m})_{m\in\mathcal{D}_{k}(\mathcal{A})}$ and $\boldsymbol{n}=(n_{m})_{m\in\mathcal{D}_{k}(\mathcal{A})}$ where $a_{m},n_{m}\geq 0$,
$\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}f_{rad}(a_{m},n_{m})n_{m}\leq\sqrt{{\gamma}M_{k}\left(\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}a_{m}n_{m}\right)}+{\gamma}M_{k}.$
where $M_{k}=|\mathcal{D}_{k}(\mathcal{A})|$.
Lemma 3 ([9], adapted).
Let $\hat{\mu}_{m}(t)=(\sum_{s<t:m_{s}=m}r_{s})/(N_{m}(t)+1)$ be the empirical average performance and $\hat{c}_{m}(t)=(\sum_{s<t:m_{s}=m}c_{s})/(N_{m}(t)+1)$ be the empirical average cost for model $m\in\mathcal{D}_{k}(\mathcal{A})$ based on $N_{m}(t)$ plays before time $t$ within the current stage $k$. Then, for every $m\in\mathcal{D}_{k}(\mathcal{A})$ and time $t\in[\tau_{k},\tau_{k+1})$, with probability $1-e^{-\Omega({\gamma})}$ (i.e., on the event $\mathcal{E}$):
$|\hat{\mu}_{m}(t)-\mu_{m}|$ $\leq 2f_{rad}(\hat{\mu}_{m}(t),N_{m}(t)+1)$ (9)
$|\hat{c}_{m}(t)-\mathbb{E}[c_{m}]|$ $\leq 2f_{rad}(\hat{c}_{m}(t),N_{m}(t)+1)$ (10)
Proof.
Follows from applying Lemma 1 to the sequence of observed performances $r_{s}$ (for $m_{s}=m$) and observed costs $c_{s}$ (for $m_{s}=m$). For a fixed model $m$, the rewards $r_{s}$ (when $m_{s}=m$) are i.i.d. samples from $\nu_{m}(\cdot)$ with mean $\mu_{m}$. Similarly, costs $c_{s}$ (when $m_{s}=m$) are effectively i.i.d. samples with mean $\mathbb{E}[c_{m}]$. Thus, the conditional expectation $\mathbb{E}[x_{i}|x_{1},\ldots,x_{i-1}]$ in Lemma 1 becomes the true mean $\mu_{m}$ (or $\mathbb{E}[c_{m}]$). The derivation is analogous to Lemma 4.3 of [9]. For instance, for performance:
$|\hat{\mu}_{m}(t)-\mu_{m}|$ $=\left|\frac{\sum_{s<t:m_{s}=m}r_{s}}{N_{m}(t)+1}-\frac{(N_{m}(t)+1)\mu_{m}}{N_{m}(t)+1}\right|$
[…]
$\leq f_{rad}(\hat{\mu}_{m}(t),N_{m}(t)+1)+\frac{\mu_{m}}{N_{m}(t)+1}$
$\leq 2f_{rad}(\hat{\mu}_{m}(t),N_{m}(t)+1)$
The argument for cost is similar due to the assumption $c_{m}\in[c_{1},c_{2}]\subseteq[0,1]$. ∎
B.2 Regret Decomposition with Time-Varying Benchmark
To analyze the regret of StageRoute (Algorithm 1) over the horizon $T$, we decompose the total regret into components corresponding to the model deployment and request routing phases.
Definition 1 (Optimal Performance within Deployed Set).
For a given stage $k$ (time interval $[\tau_{k},\tau_{k+1})$ of length $T_{k}=\tau_{k+1}-\tau_{k}$) where StageRoute (during its Model Deployment Phase) deploys the set $\mathcal{D}_{k}=\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$, $V(b,\mathcal{D}_{k})$ represents the optimal expected reward per query achievable using only models from the deployed set $\mathcal{D}_{k}$. The total optimal expected performance within this stage using $\mathcal{D}_{k}$ is $\text{OPT}_{k}=T_{k}\cdot V(b,\mathcal{D}_{k})$. Note that $\mathcal{D}_{k}$ is determined by StageRoute based on $\mathcal{M}_{\tau_{k}}$ and estimates available at $\tau_{k}$. Thus, $\mathcal{D}_{k}$ and consequently $V(b,\mathcal{D}_{k})$ (and $\text{OPT}_{k}$) are random variables, dependent on the algorithm’s choices and observations up to time $\tau_{k}$.
Definition 2 (Algorithm Performance).
Let ALGO be the total expected reward accumulated by the algorithm over the horizon $T$:
$\text{ALGO}=\sum_{t=1}^{T}r_{t}=\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}.$
Let $\text{ALGO}_{k}=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}$ be the reward accumulated by the algorithm during stage $k$.
Lemma 4 (Regret Decomposition with Time-Varying Benchmark).
The total expected regret $\mathcal{R}(T)=\text{OPT}^{*}-\mathbb{E}[\text{ALGO}]$ of StageRoute, compared against the optimal time-varying benchmark $\text{OPT}^{*}=\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})$, can be decomposed as:
$\mathcal{R}(T)=\underbrace{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\left(V(b,\mathcal{D}_{k})-r_{t}\right)\right]}_{\mathcal{R}_{\textnormal{routing}}(T)}+\underbrace{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\left(V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k})\right)\right]}_{\mathcal{R}_{\textnormal{deploy}}(T)}$
where:
  • $\mathcal{R}_{\textnormal{routing}}(T)$ is the total expected routing regret, accumulating the per-query difference between the optimal expected performance with the deployed set $V(b,\mathcal{D}_{k})$ and the realized reward $r_{t}$, summed over all queries and stages.
  • $\mathcal{R}_{\textnormal{deploy}}(T)$ is the total expected deployment regret, accumulating the per-query difference in optimal expected performance achievable with the full set of available models $V(b,\mathcal{M}_{\tau_{k}})$ versus the deployed set $V(b,\mathcal{D}_{k})$, summed over all queries and stages.
Proof.
We start with the definition of the total expected regret:
$\mathcal{R}(T)=\text{OPT}^{*}-\mathbb{E}[\text{ALGO}].$
Using the definition $\text{OPT}^{*}=\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})$ and $\text{ALGO}=\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}$:
$\mathcal{R}(T)=\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})-\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}\right].$
Note that $\mathcal{M}_{\tau_{k}}$ (the set of models available at time $\tau_{k}$) depends on the fixed model arrival times $t_{m}$ and the stage start time $\tau_{k}$. According to the system model (Section 3), $\tau_{k}=(k-1)T/K+1$ and the stage length $T_{k}=(\tau_{k+1}-\tau_{k})=T/K$ are deterministic. Consequently, the set $\mathcal{M}_{\tau_{k}}$ and the benchmark value $V(b,\mathcal{M}_{\tau_{k}})$ are deterministic for each stage $k$. The randomness in the regret decomposition arises from the algorithm’s choices, specifically the selection of $\mathcal{D}_{k}$ (which determines $V(b,\mathcal{D}_{k})$) and the subsequent routing decisions leading to the realized rewards $r_{t}$.
Since $V(b,\mathcal{M}_{\tau_{k}})$ is deterministic for each $k$ and constant for $t\in[\tau_{k},\tau_{k+1}-1)$, the sum $\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})$ is also deterministic. Thus, it can be written as $\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})\right]$. We add and subtract the term $\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{D}_{k})\right]$:
$\mathcal{R}(T)$ $=\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})\right]-\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}\right]$
$=\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{M}_{\tau_{k}})\right]-\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{D}_{k})\right]$
$\qquad+\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{D}_{k})\right]-\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}\right].$
Now, we combine terms using the linearity of expectation:
$\mathcal{R}(T)$ $=\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\left(V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k})\right)\right]$
$\qquad+\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\left(V(b,\mathcal{D}_{k})-r_{t}\right)\right].$
This expression matches the claimed decomposition, identifying the deployment regret $\mathcal{R}_{\textnormal{deploy}}(T)$ and the routing regret $\mathcal{R}_{\textnormal{routing}}(T)$ as defined in the lemma statement. Alternatively, letting $\text{OPT}_{k}=T_{k}V(b,\mathcal{D}_{k})=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}V(b,\mathcal{D}_{k})$ and $\text{ALGO}_{k}=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}$, the routing regret can be written as $\mathbb{E}[\sum_{k=1}^{K}(\text{OPT}_{k}-\text{ALGO}_{k})]$. Similarly, the deployment regret can be written as $\mathbb{E}[\sum_{k=1}^{K}T_{k}(V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k}))]$. ∎
This decomposition provides an accurate picture of the algorithm’s performance. The deployment regret $\mathcal{R}_{\textnormal{deploy}}(T)$ isolates the loss incurred specifically by StageRoute’s potentially suboptimal selection $\mathcal{D}_{k}$ (during its Model Deployment Phase) from the available set $\mathcal{M}_{\tau_{k}}$, measured against the best possible rate $V(b,\mathcal{M}_{\tau_{k}})$ achievable with those available models. Bounding $\mathcal{R}_{\textnormal{deploy}}(T)$ involves analyzing how effectively the Model Deployment Phase of StageRoute identifies the optimal subset of size at most $M_{\max}$ from $\mathcal{M}_{\tau_{k}}$ based on its estimates. The routing regret $\mathcal{R}_{\textnormal{routing}}(T)$ remains the sum of per-query differences between the optimal expected performance using the deployed models $\mathcal{D}_{k}$ and the actual realized rewards $r_{t}$. Lemma 11 (or subsequent analysis) addresses the term $\mathbb{E}[\sum_{t=\tau_{k}}^{\tau_{k+1}-1}(V(b,\mathcal{D}_{k})-r_{t})\mid\mathcal{D}_{k}]$ which contributes to $\mathcal{R}_{\textnormal{routing}}(T)$.
B.3 Analysis of Deployment Regret
We now analyze the deployment regret component $\mathcal{R}_{\textnormal{deploy}}(T)$ as defined in Lemma 4:
$\mathcal{R}_{\textnormal{deploy}}(T)=\mathbb{E}\left[\sum_{k=1}^{K}T_{k}\left(V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k})\right)\right].$
This quantity captures the cumulative expected performance loss across all stages, incurred when the StageRoute algorithm selects a subset $\mathcal{D}_{k}$ at stage $k$ based on estimated model statistics at time $\tau_{k}$, instead of deploying the optimal subset from the full set of available models $\mathcal{M}_{\tau_{k}}$.
The deployment regret arises from two complementary sources:
  1. 1.
    Parameter Uncertainty: Inaccurate estimates of model performance ($\mu_{m}$) and cost ($\mathbb{E}[c_{m}]$) may result in suboptimal deployment decisions. This source corresponds to models that have already been deployed in one or more of the previous $k-1$ stages.
  2. 2.
    Model Discovery Bottleneck: The constraint that at most $M_{\max}$ models can be deployed concurrently may exclude promising but underexplored models—particularly newly arrived ones—from being included in $\mathcal{D}_{k}$. This prevents timely evaluation and utilization, contributing to additional regret. This case pertains to models that have not been selected in any of the preceding $k-1$ stages.
These two components are complementary in nature and together constitute the total deployment regret $\mathcal{R}_{\textnormal{deploy}}(T)$. In the following analysis, we will provide regret bounds for each source.
For the analysis, we recall that $V(b,\mathcal{S})$ is the optimal performance rate for a model set $\mathcal{S}$ (defined in Section 3). Let $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ be the optimal rate achievable using all models available at the start of stage $k$, and $V_{k}=V(b,\mathcal{D}_{k})$ be the optimal rate achievable using the subset $\mathcal{D}_{k}\subseteq\mathcal{M}_{\tau_{k}}$ selected by StageRoute. $V^{*}_{k}$ is deterministic given $\tau_{k}$, while $V_{k}$ (through $\mathcal{D}_{k}$) is a random variable. A key assumption for bounding the deployment regret due to parameter uncertainty involves the Lagrange multipliers associated with the budget constraint in the definition of $V^{*}_{k}$. Let $\lambda^{*}_{k}\geq 0$ be this optimal Lagrange multiplier. We assume that $\lambda^{*}_{k}$ is uniformly bounded by a constant $\Lambda$ for all $k$. This is a common assumption in the analysis of learning algorithms with budget constraints and is often justified when standard regularity conditions (such as Slater’s condition, which we assume in Assumption 1) hold for the underlying optimization problems, particularly given that our problem parameters (rewards, costs, capacities $\alpha_{m}$) are bounded and expected costs $\mathbb{E}[c_{m}]$ are lower-bounded by $c_{1}>0$.
Confidence Bounds and Good Event.
Let $\tau_{k}$ be the start of stage $k$. $N_{m}(\tau_{k})$ is the play count for model $m$ before $\tau_{k}$. Define confidence radii using Lemma 1 and 3 with $\gamma=\Theta(\log(NT/\delta))$:
$\text{rad}_{\mu}(m,\tau_{k})$ $=2f_{rad}(\bar{\mu}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$
$\text{rad}_{c}(m,\tau_{k})$ $=2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)\quad(\text{Since costs }c_{m}\in[c_{1},c_{2}]\subseteq[0,1])$
Let $\mathcal{E}_{k}$ be the good event at time $\tau_{k}$ where, for all $m\in\mathcal{M}_{\tau_{k}}$, the confidence bounds based on $\gamma$ hold:
$\mu_{m}$ $\leq\mu^{U}_{m}(\tau_{k})\quad\text{and}\quad\mu^{U}_{m}(\tau_{k})\leq\mu_{m}+\text{rad}_{\mu}(m,\tau_{k})$
$\mathbb{E}[c_{m}]-\text{rad}_{c}(m,\tau_{k})$ $\leq c^{L}_{m}(\tau_{k})\quad\text{and}\quad c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$
Let $\mathcal{E}=\cap_{k=1}^{K}\mathcal{E}_{k}$. By a union bound over all $N$ models in the universe $\mathcal{M}_{T}$ and $K$ stages, $\mathbb{P}(\mathcal{E})\geq 1-\delta$. We condition the analysis on $\mathcal{E}$. (Note: $\mu^{U}_{m}$ is $\bar{\mu}_{m}(\tau_{k})+2f_{rad}(\bar{\mu}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$ and $c^{L}_{m}$ is $\bar{c}_{m}(\tau_{k})-2f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$ as per Eq. (6,7) in the algorithm description, projected onto $[0,1]$ and $[c_{1},c_{2}]$ respectively. The inequalities above capture the desired properties on the good event $\mathcal{E}_{k}$).
Bounding the Per-Stage Deployment Gap due to Estimation Uncertainty.
Lemma 5 (Per-Stage Deployment Gap Bound).
Let $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ be the optimal rate with available models at stage $k$. Let $\mathcal{D}_{k}=\textnormal{supp}(d^{*})$ be the set selected by the Model Deployment Phase of StageRoute based on $\mathcal{M}_{\tau_{k}}$ and estimates at $\tau_{k}$, via solution $d^{*}$ (from Eq. (3)). Let $V_{k}=V(b,\mathcal{D}_{k})$. On the good event $\mathcal{E}_{k}$, the deployment gap for stage $k$ is bounded as:
$V^{*}_{k}-V_{k}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\left(\textnormal{rad}_{\mu}(m,\tau_{k})+\lambda^{*}_{k}\textnormal{rad}_{c}(m,\tau_{k})\right)d^{*}_{m}$
where $\lambda^{*}_{k}$ is the optimal dual variable for the budget constraint in the problem defining $V^{*}_{k}$, assumed to be $\leq\Lambda$.
Proof.
Let $d^{\textnormal{opt},k}$ be an optimal solution achieving $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$. Let $d^{*}$ be the solution found by StageRoute’s Model Deployment Phase (using Eq. (3)) at $\tau_{k}$ when optimizing over $\mathcal{M}_{\tau_{k}}$ using $\mu^{U}_{m}(\tau_{k})$ and $c^{L}_{m}(\tau_{k})$. Let $\mathcal{D}_{k}=\textnormal{supp}(d^{*})$. Let $V_{k}=V(b,\mathcal{D}_{k})$.
On the event $\mathcal{E}_{k}$, the confidence bounds hold for all $m\in\mathcal{M}_{\tau_{k}}$. Specifically, $\mu_{m}\leq\mu^{U}_{m}(\tau_{k})\leq\mu_{m}+\text{rad}_{\mu}(m,\tau_{k})$ and $\mathbb{E}[c_{m}]-\text{rad}_{c}(m,\tau_{k})\leq c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$. The true optimal solution $d^{\textnormal{opt},k}$ for the set $\mathcal{M}_{\tau_{k}}$ satisfies $\sum_{m\in\mathcal{M}_{\tau_{k}}}\mathbb{E}[c_{m}]d^{\textnormal{opt},k}_{m}\leq b$, uses $\leq M_{\max}$ models from $\mathcal{M}_{\tau_{k}}$, etc. Since $c^{L}_{m}(\tau_{k})\leq\mathbb{E}[c_{m}]$ on $\mathcal{E}_{k}$, we have $\sum_{m\in\mathcal{M}_{\tau_{k}}}c^{L}_{m}(\tau_{k})d^{\textnormal{opt},k}_{m}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\mathbb{E}[c_{m}]d^{\textnormal{opt},k}_{m}\leq b$. Thus, $d^{\textnormal{opt},k}$ is a feasible solution for the optimization problem solved by StageRoute (Eq. (3) applied to $\mathcal{M}_{\tau_{k}}$).
By the optimality of $d^{*}$ for StageRoute’s deployment objective over $\mathcal{M}_{\tau_{k}}$:
$\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu^{U}_{m}(\tau_{k})d^{*}_{m}\geq\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu^{U}_{m}(\tau_{k})d^{\textnormal{opt},k}_{m}$ (11)
Using the confidence bounds on $\mathcal{E}_{k}$ for $m\in\mathcal{M}_{\tau_{k}}$:
$\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu^{U}_{m}(\tau_{k})d^{*}_{m}$ $\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}(\mu_{m}+\text{rad}_{\mu}(m,\tau_{k}))d^{*}_{m}=\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu_{m}d^{*}_{m}+\sum_{m\in\mathcal{M}_{\tau_{k}}}\text{rad}_{\mu}(m,\tau_{k})d^{*}_{m}$
$\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu^{U}_{m}(\tau_{k})d^{\textnormal{opt},k}_{m}$ $\geq\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu_{m}d^{\textnormal{opt},k}_{m}=V^{*}_{k}$
Substituting these into Eq. (11):
$\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu_{m}d^{*}_{m}+\sum_{m\in\mathcal{M}_{\tau_{k}}}\text{rad}_{\mu}(m,\tau_{k})d^{*}_{m}\geq V^{*}_{k}$
Rearranging:
$\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu_{m}d^{*}_{m}\geq V^{*}_{k}-\sum_{m\in\mathcal{M}_{\tau_{k}}}\text{rad}_{\mu}(m,\tau_{k})d^{*}_{m}$ (12)
This bounds the true performance of the distribution $d^{*}$ chosen by the algorithm. Now we relate this to $V_{k}=V(b,\mathcal{D}_{k})$, the optimal performance within the chosen set $\mathcal{D}_{k}=\textnormal{supp}(d^{*})\subseteq\mathcal{M}_{\tau_{k}}$. The distribution $d^{*}$ is supported on $\mathcal{D}_{k}$ and uses at most $M_{\max}$ models (due to the constraint in Eq. (3)). We examine its feasibility w.r.t. the true budget constraint. On $\mathcal{E}_{k}$:
$\sum_{m\in\mathcal{D}_{k}}\mathbb{E}[c_{m}]d^{*}_{m}\leq\sum_{m\in\mathcal{D}_{k}}(c^{L}_{m}(\tau_{k})+\text{rad}_{c}(m,\tau_{k}))d^{*}_{m}\leq b+\sum_{m\in\mathcal{D}_{k}}\text{rad}_{c}(m,\tau_{k})d^{*}_{m}$
Let $\delta_{c}(d^{*})=\sum_{m\in\mathcal{D}_{k}}\text{rad}_{c}(m,\tau_{k})d^{*}_{m}$. Using sensitivity analysis/duality, relating $V_{k}=V(b,\mathcal{D}_{k})$ to the performance of $d^{*}$ which is feasible for budget $b+\delta_{c}(d^{*})$:
$V_{k}=V(b,\mathcal{D}_{k})\geq V(b+\delta_{c}(d^{*}),\mathcal{D}_{k})-\lambda^{*}_{k}\delta_{c}(d^{*})$
where $\lambda^{*}_{k}\leq\Lambda$. Since $d^{*}$ is feasible for $V(b+\delta_{c}(d^{*}),\mathcal{D}_{k})$:
$V(b+\delta_{c}(d^{*}),\mathcal{D}_{k})\geq\sum_{m\in\mathcal{D}_{k}}\mu_{m}d^{*}_{m}=\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu_{m}d^{*}_{m}$
Combining these:
$V_{k}$ $\geq\left(\sum_{m\in\mathcal{M}_{\tau_{k}}}\mu_{m}d^{*}_{m}\right)-\lambda^{*}_{k}\delta_{c}(d^{*})$
[…]
Rearranging gives the result (noting $d^{*}_{m}=0$ for $m\notin\mathcal{D}_{k}$):
$V^{*}_{k}-V_{k}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\text{rad}_{\mu}(m,\tau_{k})d^{*}_{m}+\lambda^{*}_{k}\sum_{m\in\mathcal{D}_{k}}\text{rad}_{c}(m,\tau_{k})d^{*}_{m}$
Since $d^{*}_{m}=0$ for $m\notin\mathcal{D}_{k}$, the second sum can also be written over $\mathcal{M}_{\tau_{k}}$:
$V^{*}_{k}-V_{k}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\left(\text{rad}_{\mu}(m,\tau_{k})+\lambda^{*}_{k}\text{rad}_{c}(m,\tau_{k})\right)d^{*}_{m}$
Cumulative Deployment Regret from Estimation Uncertainty.
Summing the per-stage deployment gaps caused by estimation errors gives the learning component of the deployment regret.
Lemma 6 (Deployment Regret from Estimation Uncertainty).
Assume the optimal dual variables $\lambda^{*}_{k}$ are uniformly bounded by $\Lambda$. Let $K$ be the total number of stages. Set the confidence parameter $\gamma=\Theta(\log(NT/\delta))$, where $N=|\mathcal{M}_{T}|$. Then the component of total expected deployment regret due to parameter uncertainty, denoted $\mathcal{R}_{\textnormal{deploy,learn}}(T)$, is bounded by:
$\mathcal{R}_{\textnormal{deploy,learn}}(T)\leq\mathcal{O}\left(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})}+M_{\max}K\log(NT/\delta)\right).$
Proof.
The total expected deployment regret due to parameter uncertainty is given by
$\mathcal{R}_{\textnormal{deploy,learn}}(T)=\mathbb{E}\left[\sum_{k=1}^{K}T_{k}\left(V(b,\mathcal{M}_{\tau_{k}})-V(b,\mathcal{D}_{k})\right)\right],$
where $V^{*}_{k}=V(b,\mathcal{M}_{\tau_{k}})$ and $V_{k}=V(b,\mathcal{D}_{k})$. We condition on the good event $\mathcal{E}=\cap_{k=1}^{K}\mathcal{E}_{k}$, which holds with probability at least $1-\delta$. On this event, the confidence bounds for $\mu_{m}$ and $\mathbb{E}[c_{m}]$ hold for all models $m$ and stages $k$. From Lemma 5, on event $\mathcal{E}_{k}$:
$V^{*}_{k}-V_{k}\leq\sum_{m\in\mathcal{M}_{\tau_{k}}}\left(\text{rad}_{\mu}(m,\tau_{k})+\lambda^{*}_{k}\text{rad}_{c}(m,\tau_{k})\right)d^{*}_{k,m}.$
Let $C_{m}(\tau_{k})=\text{rad}_{\mu}(m,\tau_{k})+\Lambda\text{rad}_{c}(m,\tau_{k})$, using the uniform bound $\lambda^{*}_{k}\leq\Lambda$. Since $d^{*}_{k,m}=0$ for $m\notin\mathcal{D}_{k}=\text{supp}(d^{*}_{k})$, we have:
$T_{k}(V^{*}_{k}-V_{k})\leq T_{k}\sum_{m\in\mathcal{D}_{k}}C_{m}(\tau_{k})d^{*}_{k,m}.$
The term $d^{*}_{k,m}$ represents the selection weight for model $m$ in the deployment optimization at stage $k$. Thus, $\sum_{m\in\mathcal{D}_{k}}C_{m}(\tau_{k})d^{*}_{k,m}$ is a weighted average of the combined confidence radii for the deployed models. We are considering the case where the model has already been selected in the previous $k-1$ stages. Thus, we can bound the sum of $T_{k}(V^{*}_{k}-V_{k})$ by terms related to $\sum_{k=1}^{K}\sum_{m\in\mathcal{D}_{k}}n_{k,m}C_{m}(\tau_{k})$. Specifically,
$\sum_{k=1}^{K}T_{k}(V^{*}_{k}-V_{k})\leq\mathcal{O}(1)\sum_{k=1}^{K}\sum_{m\in\mathcal{D}_{k}}n_{k,m}C_{m}(\tau_{k}).$
Let’s analyze the sum $S=\sum_{k=1}^{K}\sum_{m\in\mathcal{D}_{k}}n_{k,m}C_{m}(\tau_{k})$. Recall $C_{m}(\tau_{k})=2f_{rad}(\bar{\mu}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)+2\Lambda f_{rad}(\bar{c}_{m}(\tau_{k}),N_{m}(\tau_{k})+1)$. Since rewards $\bar{\mu}_{m}(\tau_{k})\in[0,1]$ and costs $\bar{c}_{m}(\tau_{k})\in[c_{1},c_{2}]\subseteq[0,1]$ (with $c_{1}>0$), we have $f_{rad}(v,n)=\sqrt{\frac{\gamma v}{n}}+\frac{\gamma}{n}\leq\sqrt{\frac{\gamma}{n}}+\frac{\gamma}{n}$ for $v\in[0,1]$. So, $C_{m}(\tau_{k})\leq 2(1+\Lambda)\left(\sqrt{\frac{\gamma}{N_{m}(\tau_{k})+1}}+\frac{\gamma}{N_{m}(\tau_{k})+1}\right)$. Let $C^{\prime}=2(1+\Lambda)$.
$S\leq C^{\prime}\sum_{k=1}^{K}\sum_{m\in\mathcal{D}_{k}}n_{k,m}\left(\sqrt{\frac{\gamma}{N_{m}(\tau_{k})+1}}+\frac{\gamma}{N_{m}(\tau_{k})+1}\right).$
We can rewrite the sum by first summing over all models $m\in\mathcal{M}_{T}$ (the set of all $N$ possible models) and then over the stages $k$ in which $m$ was deployed and played:
$S\leq C^{\prime}\sum_{m\in\mathcal{M}_{T}}\sum_{k:m\in\mathcal{D}_{k}\text{ and }n_{k,m}>0}n_{k,m}\left(\sqrt{\frac{\gamma}{N_{m}(\tau_{k})+1}}+\frac{\gamma}{N_{m}(\tau_{k})+1}\right).$
For a fixed model $m$, let $N_{m}(T)$ be the total number of times it is played up to $T$. Let $n_{m,s_{j}}$ be the number of times $m$ is played during the $j$-th stage (denoted $s_{j}$) in which it is deployed. Let $N_{m}(\tau_{s_{j}})$ be the total number of plays of $m$ before stage $s_{j}$. The sum for model $m$ is:
$S_{m}=\sum_{j=1}^{K^{\prime}_{m}}n_{m,s_{j}}\left(\sqrt{\frac{\gamma}{N_{m}(\tau_{s_{j}})+1}}+\frac{\gamma}{N_{m}(\tau_{s_{j}})+1}\right),$
where $K^{\prime}_{m}$ is the number of stages model $m$ is played. We use the standard inequalities for such sums:
  • $\sum_{j=1}^{K^{\prime}_{m}}n_{m,s_{j}}\sqrt{\frac{\gamma}{N_{m}(\tau_{s_{j}})+1}}\leq\sqrt{\gamma}\sum_{i=1}^{N_{m}(T)}\frac{1}{\sqrt{(\text{plays of }m\text{ before current block})+1}}\leq\sqrt{\gamma}\cdot 2\sqrt{N_{m}(T)}$.
  • $\sum_{j=1}^{K^{\prime}_{m}}n_{m,s_{j}}\frac{\gamma}{N_{m}(\tau_{s_{j}})+1}\leq\gamma\sum_{i=1}^{N_{m}(T)}\frac{1}{(\text{plays of }m\text{ before current block})+1}\leq\gamma\cdot(1+\ln N_{m}(T))$.
So, $S_{m}\leq\mathcal{O}(\sqrt{\gamma N_{m}(T)}+\gamma\log N_{m}(T))$. (We absorb constants into $\mathcal{O}(\cdot)$ for now, and will re-introduce $C^{\prime}$ later). Thus, $S\leq C^{\prime}\sum_{m\in\mathcal{M}_{T}\text{ s.t. }N_{m}(T)>0}\mathcal{O}(\sqrt{\gamma N_{m}(T)}+\gamma\log N_{m}(T))$.
The sum is over models that were actually played. Let $\mathcal{M}_{P}=\{m\in\mathcal{M}_{T}\mid N_{m}(T)>0\}$. The number of models in $\mathcal{D}_{k}$ is $|\mathcal{D}_{k}|\leq M_{\max}$. So, $\sum_{m\in\mathcal{M}_{P}}\sqrt{N_{m}(T)}\leq\sqrt{|\mathcal{M}_{P}|\sum_{m\in\mathcal{M}_{P}}N_{m}(T)}=\sqrt{|\mathcal{M}_{P}|T}$ by Cauchy-Schwarz. Since at most $M_{\max}$ models are deployed in any stage, and there are $K$ stages, the number of distinct models ever deployed is $|\mathcal{M}_{P}|\leq\min(N,KM_{\max})$.
The sum $\sum_{m\in\mathcal{M}_{P}}\log N_{m}(T)$. If $N_{m}(T)\geq 1$, then $\log N_{m}(T)\geq 0$. This sum is at most $M_{\max}\log(T/M_{\max})$ if $M_{\max}$ models share $T$ plays, or more generally bounded by $M_{\max}K$ (if each of $M_{\max}$ models gets played at least once in each of $K$ stages, its $\log N_{m}(T)$ contributes, and $N_{m}(T)$ could be small). A more careful bound for the sum of $\log$ terms: $\sum_{m\in\mathcal{M}_{P}}\log N_{m}(T)\leq M_{\max}K\log T$. So, $S\leq C^{\prime}\left(\mathcal{O}(\sqrt{\gamma T\cdot\min(N,KM_{\max})})+\mathcal{O}(\gamma M_{\max}K)\right)$. The $\log$ terms are typically absorbed into the $\gamma$ term. The expectation $\mathbb{E}[S]$ includes the good event (probability $1-\delta$) and the bad event (probability $\delta$). On the bad event, the regret in one stage is at most $T_{k}$, so total $T$. $\mathcal{R}_{\textnormal{deploy,learn}}(T)\leq\mathbb{E}[S]+\delta T$. If $\delta=\mathcal{O}(1/T)$, then $\delta T=\mathcal{O}(1)$. Substituting $C^{\prime}=2(1+\Lambda)$ and $\gamma=\Theta(\log(NT/\delta))$, then we complete the proof. ∎
Deployment Regret from Model Discovery Bottleneck.
The constraint of deploying at most $M_{\max}$ models simultaneously, $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$, introduces a structural challenge, particularly when new models frequently become available or the total pool of models $\mathcal{M}_{T}$ is large. This challenge is the model discovery bottleneck: identifying truly superior models among many new, unevaluated candidates can be delayed.
The StageRoute algorithm employs UCBs for rewards ($\mu^{U}_{m}$) and LCBs for costs ($c^{L}_{m}$) in its DeployOPT phase (Eq. (3)). For a model $m$ that is newly available at stage $k$ (i.e., $t_{m}\leq\tau_{k}$ and its count of previous selections $N_{m}(\tau_{k})=0$), its initial empirical averages $\bar{\mu}_{m}(\tau_{k})$ and $\bar{c}_{m}(\tau_{k})$ are set based on priors. The confidence radius $f_{rad}(v,N_{m}(\tau_{k})+1)=\sqrt{\frac{\gamma v}{N_{m}(\tau_{k})+1}}+\frac{\gamma}{N_{m}(\tau_{k})+1}$ becomes large for $N_{m}(\tau_{k})=0$. Specifically, with $N_{m}(\tau_{k})+1=1$:
$\mu^{U}_{m}(\tau_{k})$ $=\textnormal{proj}_{[0,1]}\left(\bar{\mu}_{m}(\tau_{k})+2\left(\sqrt{\gamma\bar{\mu}_{m}(\tau_{k})}+\gamma\right)\right),$
$c^{L}_{m}(\tau_{k})$ $=\textnormal{proj}_{[c_{1},c_{2}]}\left(\bar{c}_{m}(\tau_{k})-2\left(\sqrt{\gamma\bar{c}_{m}(\tau_{k})}+\gamma\right)\right).$
Assuming priors are chosen such that new models are treated optimistically (or if $\gamma$ is sufficiently large), $\mu^{U}_{m}(\tau_{k})$ will be close to 1 (e.g., if $\bar{\mu}_{m}(\tau_{k})=0$, $\mu^{U}_{m}(\tau_{k})=\textnormal{proj}_{[0,1]}(2\gamma)\approx 1$ for appropriate $\gamma$) and $c^{L}_{m}(\tau_{k})$ will be close to $c_{1}$ (e.g., if $\bar{c}_{m}(\tau_{k})=c_{1}$, $c^{L}_{m}(\tau_{k})=\textnormal{proj}_{[c_{1},c_{2}]}(c_{1}-2(\sqrt{\gamma c_{1}}+\gamma))\approx c_{1}$, noting $c_{1},c_{2}\in[0,1]$). Let these optimistic initial values be $U_{init}$ and $L_{init}$ respectively.
Consider an update point $\tau_{k}$. Let $\mathcal{N}_{new,k}\subseteq\mathcal{M}_{\tau_{k}}$ be the set of models that are new at or before $\tau_{k}$ and have not yet been deployed ($N_{m}(\tau_{k})=0$). All models in $\mathcal{N}_{new,k}$ will have nearly identical, highly optimistic $(\mu^{U}_{m},c^{L}_{m})\approx(U_{init},L_{init})$ values. If the number of such equally optimistic new models, $|\mathcal{N}_{new,k}|$, plus other potentially optimistic (but previously explored) models, exceeds $M_{\max}$, the DeployOPT phase must select only $M_{\max}$ models. If the new models in $\mathcal{N}_{new,k}$ dominate the selection pool due to their optimism, DeployOPT will choose $M_{\max}$ models from $\mathcal{N}_{new,k}$ (possibly along with some already explored models). Crucially, if there are more than $M_{\max}$ models within $\mathcal{N}_{new,k}$ (or a larger pool of similarly optimistic candidates) that yield effectively the same objective value for DeployOPT (because their $\mu^{U}_{m},c^{L}_{m},\alpha_{m}$ are similar), the selection among these specific candidates becomes arbitrary (e.g., dependent on tie-breaking rules).
A truly superior new model $m^{*}\in\mathcal{N}_{new,k}$ might thus be part of a large batch of $N_{batch}>M_{\max}$ new models that all appear equally promising to DeployOPT. In this scenario, $m^{*}$ might not be selected for deployment in stage $k$, deferring its evaluation. This deferral means the system misses the opportunity to benefit from $m^{*}$’s potentially high true performance $\mu_{m^{*}}$ for the duration of stage $k$, which is $T_{k}=T/K$ rounds.
Lemma 7 (Model Discovery Bottleneck Regret).
Let $M_{\max}$ be the maximum number of concurrently deployed models, $T$ be the total time horizon, and $K$ be the number of stages, with each stage having $T_{k}=T/K$ rounds. The component of expected deployment regret due to the bottleneck in discovering and evaluating all $N$ models, denoted $\mathcal{R}_{\textnormal{deploy,discovery}}(T)$, is bounded by:
$\mathcal{R}_{\textnormal{deploy,discovery}}(T)=\mathcal{O}\left(\frac{N\cdot(T/K)}{M_{\max}}\right)=\mathcal{O}\left(\frac{NT}{M_{\max}K}\right).$
Proof.
Each of the $N$ models in the universe $\mathcal{M}_{T}$ needs to be deployed at least once to gather initial empirical data and move its UCB/LCB estimates away from their initial purely optimistic values. We are interested in the total regret incurred until all $N$ models have had at least one such initial deployment opportunity.
Due to the constraint $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$, at most $M_{\max}$ distinct models can be deployed and evaluated in any given stage $k$. If the system prioritizes exploring previously undeployed models (which is encouraged by their optimistic UCB/LCB values), it will take a minimum of $K_{explore}=\lceil N/M_{\max}\rceil$ stages to ensure that every model in $\mathcal{M}_{T}$ (assuming all become available early enough) has been deployed at least once.
Consider a discovery period spanning these first $K_{explore}$ effective stages. During any given stage $j\in[1,K_{explore}]$ within this period, if the set of $M_{\max}$ models deployed, $\mathcal{D}_{j}(\mathcal{A})$, does not include some model $m^{*}\in\mathcal{M}_{T}$ which is (a) available ($m^{*}\in\mathcal{M}_{\tau_{j}}$), (b) truly superior to at least one deployed model $m^{\prime}\in\mathcal{D}_{j}(\mathcal{A})$, and (c) $m^{*}$ has not been deployed yet because it is waiting its turn due to the arbitrary selection among many new, equally optimistic models, then regret is incurred. The per-query regret in such a case can be up to $1$ (if $\mu_{m^{*}}\approx 1$ and $\mu_{m^{\prime}}\approx 0$).
The total number of queries over these $K_{explore}$ stages is $K_{explore}\cdot T_{k}=\lceil N/M_{\max}\rceil\cdot(T/K)$. During this cumulative period, the system is effectively cycling through the $N$ models. If the selection process within each stage means that, on average, the deployed set is suboptimal because truly good models are among the $(N-j\cdot M_{\max})$ yet to be tried, the system incurs regret. The term $\mathcal{O}(NT/(M_{\max}K))$ represents the cumulative regret if, for a duration equivalent to $N/M_{\max}$ full stages (each of length $T/K$), the system operates with a deployed set that is, on average, $\mathcal{O}(1)$ worse per query than if all models had already been evaluated. This occurs because the $M_{\max}$ slots are occupied by models chosen optimistically, and a superior model might be consistently deferred if it’s part of a large pool of indistinguishably optimistic new models.
More formally, consider the $N$ models. Each requires roughly one exploration slot of duration $T_{k}$. These $N$ slots are processed in parallel groups of $M_{\max}$. This implies approximately $N/M_{\max}$ stages are spent ensuring all models receive initial evaluation. If, during these $N/M_{\max}$ stages, the average deployed set yields $\mathcal{O}(1)$ less reward per query compared to an optimal deployment (had all models been known), the total regret from this discovery phase is $(N/M_{\max})\cdot(T/K)\cdot\mathcal{O}(1)$. The expectation $\mathbb{E}[\cdot]$ in $\mathcal{R}_{\textnormal{deploy,discovery}}(T)$ averages over the random tie-breaking in DeployOPT when faced with multiple equally optimistic new models, and the stochastic arrival pattern of models. The $\mathcal{O}(\cdot)$ notation absorbs constants related to the maximum possible per-query regret (e.g., 1) and the precise nature of average suboptimality during this discovery period. ∎
This component accounts for the scenarios where truly good models might be systematically delayed in their initial deployment if they frequently arrive alongside many other new models, leading to arbitrary choices among a large pool of initially indistinguishable (optimistic) candidates, subject to the $M_{\max}$ deployment limit over $K$ stages.
Total Deployment Regret.
The total deployment regret $\mathcal{R}_{\textnormal{deploy}}(T)$ is the sum of the regret from parameter uncertainty (Lemma 6) and the regret from the model discovery bottleneck (Lemma 7).
Lemma 8 (Total Deployment Regret Bound).
Let $N$ be the total number of models, $M_{\max}$ the maximum deployed models, $T$ the horizon, and $K$ the number of stages. Let $\gamma=\Theta(\log(NT/\delta))$. The total expected deployment regret $\mathcal{R}_{\textnormal{deploy}}(T)$ is bounded by:
$\mathcal{R}_{\textnormal{deploy}}(T)\leq\mathcal{O}\left(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})}+M_{\max}K\log(NT/\delta)+\frac{NT}{M_{\max}K}\right).$
Proof.
The total deployment regret is the sum of the bounds from Lemma 6 and Lemma 7:
$\mathcal{R}_{\textnormal{deploy}}(T)$ $=\mathcal{R}_{\textnormal{deploy,learn}}(T)+\mathcal{R}_{\textnormal{deploy,discovery}}(T)$
$\leq\mathcal{O}\left(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})}+M_{\max}K\log(NT/\delta)\right)+\mathcal{O}\left(\frac{NT}{M_{\max}K}\right)$
$=\mathcal{O}\left(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})}+M_{\max}K\log(NT/\delta)+\frac{NT}{M_{\max}K}\right).$
The total deployment regret bound in Lemma 8 highlights two distinct challenges in the deployment phase. The terms $\mathcal{O}(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})})$ and $\mathcal{O}(M_{\max}K\log(NT/\delta))$ capture the cost of learning the parameters of models that are considered for deployment. This cost depends on the horizon $T$, the number of deployment slots $M_{\max}$, the number of stages $K$, and logarithmic factors related to the total number of models $N$ and confidence $\delta$. The term $\mathcal{O}(NT/(M_{\max}K))$ reflects the structural cost imposed by the discovery bottleneck: when the universe of models $N$ is large compared to the deployment capacity $M_{\max}$ and the number of adaptation opportunities $K$, there is an inherent regret incurred in sequentially exploring models to identify the best ones. This term can dominate if $N$ is very large or $K$ is small, underscoring the importance of the deployment frequency and capacity in dynamic LLM environments.
B.4 Bounding the Total Routing Regret
We now bound the total routing regret term $\mathcal{R}_{\textnormal{routing}}(T)$ identified in Lemma 4. This involves summing the per-stage regrets incurred by the Request Routing Phase of StageRoute due to using estimated model parameters within the deployed sets $\mathcal{D}_{k}$. This part of the proof follows a similar structure to the analysis of UCB-style algorithms for the Bandits with Knapsacks problem [5].
Lemma 9 (Performance Bound).
Let $\text{ALGO}_{k}=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}$ be the total observed performance in stage $k$. With probability at least $1-(M_{k}T_{k})\exp(-\mathcal{O}({\gamma}))$,
$\left|\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\left(r_{t}-\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu^{U}_{m}(t)p_{t}^{*}(m)\right)\right|\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}\text{ALGO}_{k}}+{\gamma}M_{k}\right).$
Proof.
The proof follows the structure of Lemma 4.4 in [9], adapted to our notation for models $m\in\mathcal{D}_{k}(\mathcal{A})$ and performance $\mu_{m}$, UCB $\mu^{U}_{m}(t)$, chosen model $m_{t}$, routing distribution $p_{t}^{*}$ (from Eq. (7)), and summing over $t\in[\tau_{k},\tau_{k+1}-1]$ (length $T_{k}$).
We use Lemma 1 and Lemma 3 (for performance). High probability bounds analogous to (5) and (6) in the source proof hold for sums over $t\in[\tau_{k},\tau_{k+1}-1]$:
$\left|\sum_{t}(r_{t}-\mu_{m_{t}})\right|$ $\leq\mathcal{O}(T_{k}\cdot f_{rad}(\frac{1}{T_{k}}\sum_{t}\mu_{m_{t}},T_{k}))$
$\left|\sum_{t}\left(\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu^{U}_{m}(t)p_{t}^{*}(m)-\mu^{U}_{m_{t}}(t)\right)\right|$ $\leq\mathcal{O}(T_{k}\cdot f_{rad}(\frac{1}{T_{k}}\sum_{t}\mu^{U}_{m_{t}}(t),T_{k}))$
And analogous to (7) in the source, using Lemma 3 and Lemma 2:
$\left|\sum_{t=\tau_{k}}^{\tau_{k+1}-1}(\mu_{m_{t}}-\mu^{U}_{m_{t}}(t))\right|$ $\leq\sum_{t=\tau_{k}}^{\tau_{k+1}-1}|\mu_{m_{t}}-\mu^{U}_{m}(t)|$
$\leq\mathcal{O}\left(\sum_{t=\tau_{k}}^{\tau_{k+1}-1}f_{rad}(\hat{\mu}_{m_{t}}(t),N_{m_{t}}(t)+1)\right)$
$\leq\mathcal{O}\left(\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\sum_{\text{plays }j\text{ of }m\text{ in stage }k}f_{rad}(\hat{\mu}_{m}(\text{at play }j),N_{m}(\text{at play }j)+1)\right)$
$\leq\mathcal{O}\left(\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\left(\sqrt{\gamma(N_{m}(\tau_{k+1})-N_{m}(\tau_{k}))\mu_{m}}+\gamma\right)\right)$
$\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}\left(\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu_{m}(N_{m}(\tau_{k+1})-N_{m}(\tau_{k}))\right)}+{\gamma}M_{k}\right)$
$\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}\left(\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\mu_{m_{t}}\right)}+{\gamma}M_{k}\right)$
Let $A=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu^{U}_{m}(t)p_{t}^{*}(m)$. Combining these bounds using the triangle inequality on $r_{t}-\sum_{m}\mu^{U}_{m}(t)p_{t}^{*}(m)=(r_{t}-\mu_{m_{t}})+(\mu_{m_{t}}-\mu^{U}_{m_{t}}(t))+(\mu^{U}_{m_{t}}(t)-\sum_{m}\mu^{U}_{m}(t)p_{t}^{*}(m))$, similar to the source proof structure, leads to an inequality relating $A$ and $\sum r_{t}=\text{ALGO}_{k}$. If $\sum r_{t}\approx\sum\mu_{m_{t}}$ and $A\approx\sum\mu^{U}_{m_{t}}(t)$, then the difference $|\sum(\mu_{m_{t}}-\mu^{U}_{m_{t}}(t))|$ dominates. This typically leads to $A-\mathcal{O}(\sqrt{{\gamma}M_{k}A}+{\gamma}M_{k})\leq\text{ALGO}_{k}$ (assuming $\mu^{U}$ are UCBs and $A\approx\sum\mu^{U}_{m_{t}}(t)$). This implies $\sqrt{A}\leq\sqrt{\text{ALGO}_{k}}+\mathcal{O}(\sqrt{{\gamma}M_{k}})$. Substituting this back into the bounds for the difference $|\text{ALGO}_{k}-A|$ yields the claimed result.
$\left|\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}-\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu^{U}_{m}(t)p_{t}^{*}(m)\right|\leq\mathcal{O}(\sqrt{{\gamma}M_{k}\text{ALGO}_{k}}+{\gamma}M_{k}).$
Lemma 10 (Cost Bound).
Let $\sum_{t=\tau_{k}}^{\tau_{k+1}-1}c_{t}$ be the total observed cost in stage $k$. Let $B_{k}=b\cdot T_{k}$ be the effective expected cost budget for the stage. With probability at least $1-(M_{k}T_{k})\exp(-\mathcal{O}({\gamma}))$ (i.e., on event $\mathcal{E}$),
$\left|\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\left(c_{t}-\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}c^{L}_{m}(t)p_{t}^{*}(m)\right)\right|\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}B_{k}}+{\gamma}M_{k}\right).$
Proof.
The proof mirrors that of Lemma 9 (and Lemma 4.5 in [9]), replacing performance with cost, $\mu_{m}$ with $\mathbb{E}[c_{m}]$, $\mu^{U}_{m}(t)$ with $c^{L}_{m}(t)$, and $r_{t}$ with $c_{t}$. The probability distribution $p_{t}^{*}$ is from Eq. (7). Key steps involve bounding $|\sum c_{t}-\sum\mathbb{E}[c_{m_{t}}]|$, $|\sum(\sum_{m}c^{L}_{m}(t)p_{t}^{*}(m))-\sum c^{L}_{m_{t}}(t)|$, and $|\sum\mathbb{E}[c_{m_{t}}]-\sum c^{L}_{m_{t}}(t)|$. The last term is bounded similarly using Lemma 3 and Lemma 2:
$\left|\sum_{t=\tau_{k}}^{\tau_{k+1}-1}(\mathbb{E}[c_{m_{t}}]-c^{L}_{m_{t}}(t))\right|\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}\left(\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\mathbb{E}[c_{m_{t}}]\right)}+{\gamma}M_{k}\right).$
Let $A^{\prime}=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\mathbb{E}[c_{m_{t}}]$. The algorithm ensures $\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}c^{L}_{m}(t)p_{t}^{*}(m)\leq b$ at each step $t$. Summing over the stage gives $\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}c^{L}_{m}(t)p_{t}^{*}(m)\leq b\cdot T_{k}=B_{k}$. Let $X_{c}=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}c^{L}_{m}(t)p_{t}^{*}(m)$. Then $\left|\sum_{t=\tau_{k}}^{\tau_{k+1}-1}c_{t}-X_{c}\right|\leq\mathcal{O}(\sqrt{{\gamma}M_{k}A^{\prime}}+{\gamma}M_{k})$ where $A^{\prime}$ is the sum of true expected costs. On event $\mathcal{E}$, $c^{L}_{m}(t)\leq\mathbb{E}[c_{m}]$, so $X_{c}\leq A^{\prime}$. Also, $A^{\prime}\leq X_{c}+\mathcal{O}(\sqrt{{\gamma}M_{k}A^{\prime}}+{\gamma}M_{k})$. Since $X_{c}\leq B_{k}$, it follows that $A^{\prime}\leq B_{k}+\mathcal{O}(\sqrt{{\gamma}M_{k}A^{\prime}}+{\gamma}M_{k})$. This implies $\sqrt{A^{\prime}}\leq\sqrt{B_{k}}+\mathcal{O}(\sqrt{{\gamma}M_{k}})$. Substituting this back into the concentration bounds for $|\sum c_{t}-X_{c}|$ (and noting $A^{\prime}\approx B_{k}$ in the error term’s leading order) yields the final result. ∎
Lemma 11 (Per-Stage Routing Regret Bound).
Let $\text{OPT}_{k}^{val}=\max_{p}\{\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu_{m}p(m)\mid\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mathbb{E}[c_{m}]p(m)\leq b,\sum p(m)=1,0\leq p(m)\leq\alpha_{m}\}$ be the optimal expected reward rate within stage $k$ using $\mathcal{D}_{k}(\mathcal{A})$ and true parameters. Let $\text{OPT}_{k}=T_{k}\cdot\text{OPT}_{k}^{val}$ be the total optimal expected performance in stage $k$. Let $\text{ALGO}_{k}=\sum_{t=\tau_{k}}^{\tau_{k+1}-1}r_{t}$. Assume $M_{k}\gamma\leq\mathcal{O}(B_{k})$. Then, on the event $\mathcal{E}$ (implying high probability for the bounds within the stage), the routing regret for stage $k$, conditioned on $\mathcal{D}_{k}$, is bounded by:
$\mathbb{E}[\text{OPT}_{k}-\text{ALGO}_{k}\mid\mathcal{D}_{k}]\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}\text{OPT}_{k}}+{\gamma}M_{k}\right)$
(The expectation is conditioned on the choice of $\mathcal{D}_{k}$, which itself depends on information up to $\tau_{k}$).
Proof.
On event $\mathcal{E}$, the following hold:
  1. 1.
    $\mu^{U}_{m}(t)\geq\mu_{m}$ and $c^{L}_{m}(t)\leq\mathbb{E}[c_{m}]$ for all relevant $m,t$.
  2. 2.
    Lemma 9: $|\sum_{t}(r_{t}-\sum_{m}\mu^{U}_{m}(t)p_{t}^{*}(m))|\leq\mathcal{O}(\sqrt{{\gamma}M_{k}\text{ALGO}_{k}}+{\gamma}M_{k})$.
  3. 3.
    Lemma 10: $|\sum_{t}(c_{t}-\sum_{m}c^{L}_{m}(t)p_{t}^{*}(m))|\leq\mathcal{O}(\sqrt{{\gamma}M_{k}B_{k}}+{\gamma}M_{k})$. Also, $\sum_{t}c_{t}\leq B_{k}+\mathcal{O}(\sqrt{{\gamma}M_{k}B_{k}}+{\gamma}M_{k})$.
From the algorithm’s choice of $p_{t}^{*}$ (solving Eq. (7)) and property (1):
$\sum_{t=\tau_{k}}^{\tau_{k+1}-1}\sum_{m\in\mathcal{D}_{k}(\mathcal{A})}\mu^{U}_{m}(t)p_{t}^{*}(m)\geq\text{OPT}_{k}$
Combining this with property (2):
$\text{ALGO}_{k}=\sum_{t}r_{t}\geq\sum_{t}\sum_{m}\mu^{U}_{m}(t)p_{t}^{*}(m)-\mathcal{O}(\sqrt{{\gamma}M_{k}\text{ALGO}_{k}}+{\gamma}M_{k})$
$\text{ALGO}_{k}\geq\text{OPT}_{k}-\mathcal{O}(\sqrt{{\gamma}M_{k}\text{ALGO}_{k}}+{\gamma}M_{k})$
Rearranging and assuming $\text{ALGO}_{k}\leq\text{OPT}_{k}$ (regret is non-negative):
$\text{OPT}_{k}-\text{ALGO}_{k}\leq\mathcal{O}(\sqrt{{\gamma}M_{k}\text{ALGO}_{k}}+{\gamma}M_{k})$
If $\text{ALGO}_{k}\leq\text{OPT}_{k}$, then $\sqrt{\text{ALGO}_{k}}\leq\sqrt{\text{OPT}_{k}}$.
$\text{OPT}_{k}-\text{ALGO}_{k}\leq\mathcal{O}(\sqrt{{\gamma}M_{k}\text{OPT}_{k}}+{\gamma}M_{k})$
Taking expectation conditioned on $\mathcal{D}_{k}$ (and implicitly on $\mathcal{E}$ for the bounds to hold), the result follows. The assumption $M_{k}\gamma\leq\mathcal{O}(B_{k})$ is used in concentration bounds for costs. ∎
Lemma 12 (Total Routing Regret Bound).
Let $K$ be the total number of stages. Let $M_{k}=|\mathcal{D}_{k}|\leq M_{\max}$. Let $\delta\in(0,1)$ be the desired overall confidence. Set $\gamma=\Theta(\log(NT/\delta))$. Then the total expected routing regret is bounded by:
$\mathcal{R}_{\textnormal{routing}}(T)=\mathbb{E}\left[\sum_{k=1}^{K}(\text{OPT}_{k}-\text{ALGO}_{k})\right]\leq\mathcal{O}\left(\sqrt{\gamma M_{\max}KT}+K\gamma M_{\max}\right)$
Substituting $\gamma=\Theta(\log(NT/\delta))$, this becomes:
$\mathcal{R}_{\textnormal{routing}}(T)\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+KM_{\max}\log(NT/\delta)\right)$
Proof.
From Lemma 4, $\mathcal{R}_{\textnormal{routing}}(T)=\sum_{k=1}^{K}\mathbb{E}[\text{OPT}_{k}-\text{ALGO}_{k}]$. Here $\text{OPT}_{k}=T_{k}V(b,\mathcal{D}_{k})$ and $\text{ALGO}_{k}$ is the algorithm’s reward in stage $k$. The outer expectation $\mathbb{E}[\cdot]$ averages over all randomness, including $\mathcal{D}_{k}$ and the failure of event $\mathcal{E}$ (which occurs with probability $\leq\delta$). Using law of total expectation: $\mathbb{E}[\text{OPT}_{k}-\text{ALGO}_{k}]=\mathbb{E}[\mathbb{E}[\text{OPT}_{k}-\text{ALGO}_{k}\mid\mathcal{D}_{k}]]$.
We apply Lemma 11. Choosing $\gamma=\Theta(\log(NT/\delta))$ ensures that event $\mathcal{E}$ holds with probability at least $1-\delta$. On $\mathcal{E}$:
$\mathbb{E}[\text{OPT}_{k}-\text{ALGO}_{k}\mid\mathcal{D}_{k}]\leq\mathcal{O}\left(\sqrt{{\gamma}M_{k}\text{OPT}_{k}}+{\gamma}M_{k}\right)$
Taking expectation over $\mathcal{D}_{k}$:
$\mathbb{E}[\text{OPT}_{k}-\text{ALGO}_{k}]\leq\mathcal{O}\left(\mathbb{E}\left[\sqrt{{\gamma}M_{k}\text{OPT}_{k}}\right]+{\gamma}\mathbb{E}[M_{k}]\right)+\delta\cdot T_{k}$
Summing this bound over all $K$ stages:
$\mathcal{R}_{\textnormal{routing}}(T)\leq\sum_{k=1}^{K}\mathcal{O}\left(\mathbb{E}\left[\sqrt{{\gamma}M_{k}\text{OPT}_{k}}\right]+{\gamma}\mathbb{E}[M_{k}]\right)+\delta T$
Using linearity of expectation, $M_{k}\leq M_{\max}$, Jensen’s inequality ($\mathbb{E}[\sqrt{X}]\leq\sqrt{\mathbb{E}[X]}$), and $\mathbb{E}[M_{k}]\leq M_{\max}$:
$\mathcal{R}_{\textnormal{routing}}(T)$ $\leq\mathcal{O}\left(\sum_{k=1}^{K}\mathbb{E}\left[\sqrt{{\gamma}M_{\max}\text{OPT}_{k}}\right]+\sum_{k=1}^{K}{\gamma}M_{\max}\right)+\delta T$
$\leq\mathcal{O}\left(\sqrt{\gamma M_{\max}}\sum_{k=1}^{K}\sqrt{\mathbb{E}[\text{OPT}_{k}]}+K\gamma M_{\max}\right)+\delta T$
Applying Cauchy-Schwarz: $(\sum_{k=1}^{K}\sqrt{\mathbb{E}[\text{OPT}_{k}]})^{2}\leq K\sum_{k=1}^{K}\mathbb{E}[\text{OPT}_{k}]$. Thus, $\sum_{k=1}^{K}\sqrt{\mathbb{E}[\text{OPT}_{k}]}\leq\sqrt{K\sum_{k=1}^{K}\mathbb{E}[\text{OPT}_{k}]}$. Since rewards are in $[0,1]$, $V(b,\mathcal{D}_{k})\leq 1$, so $\text{OPT}_{k}=T_{k}V(b,\mathcal{D}_{k})\leq T_{k}$. Summing over $k$: $\sum_{k=1}^{K}\text{OPT}_{k}\leq\sum_{k=1}^{K}T_{k}=T$. So, $\sum_{k=1}^{K}\mathbb{E}[\text{OPT}_{k}]\leq T$. Substituting this upper bound:
$\mathcal{R}_{\textnormal{routing}}(T)\leq\mathcal{O}\left(\sqrt{\gamma M_{\max}KT}+K\gamma M_{\max}\right)+\delta T$
If $\delta$ is chosen small enough, the $\delta T$ term is absorbed. This establishes the first form of the bound. Substituting $\gamma=\Theta(\log(NT/\delta))$ yields the second form. ∎
B.5 Total Regret Bound
The overall regret of StageRoute accounts for several sources of suboptimality. The total expected regret $\mathcal{R}(T)$ can now be understood as the sum of two main components:
  1. 1.
    $\mathcal{R}_{\textnormal{routing}}(T)$: The routing regret within deployed sets (Lemma 12).
  2. 2.
    $\mathcal{R}_{\textnormal{deploy}}(T)$: The total deployment regret, encompassing both learning uncertainty and model discovery bottleneck (Lemma 8).
Summing these bounds:
$\mathcal{R}(T)$ $=\mathcal{R}_{\textnormal{routing}}(T)+\mathcal{R}_{\textnormal{deploy}}(T)$
$\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+KM_{\max}\log(NT/\delta)\right)$
$\quad+\mathcal{O}\left(\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})}+M_{\max}K\log(NT/\delta)+\frac{NT}{M_{\max}K}\right)$
Combining terms, and noting that for $K\geq 1$, $\sqrt{M_{\max}KT\log(NT/\delta)}$ dominates or is equivalent to $\sqrt{T\log(NT/\delta)\cdot\min(N,KM_{\max})}$ and $KM_{\max}\log(NT/\delta)$:
$\mathcal{R}(T)\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+\frac{NT}{M_{\max}K}\right)$
This composite regret bound highlights different facets of the online learning problem. The first two terms capture the cost of learning model parameters (both for routing and deployment) and adapting the deployment strategy within a limited number of concurrently active models ($M_{\max}$) across $K$ stages. The third term, $\mathcal{O}(NT/(M_{\max}K))$, explicitly incorporates the penalty associated with the breadth of the model universe ($N$) relative to deployment capacity ($M_{\max}$) and the granularity of adaptation opportunities ($K$ stages). This term becomes significant if many models need to be evaluated through a narrow deployment bottleneck over a limited number of stages.
Appendix C Lower Bound for Online LLM Routing
We establish a lower bound on the regret for any online LLM routing policy. The construction considers a scenario where the set of competitive models and their performances can evolve over time, divided into batches. Within each batch, the algorithm faces a sequential decision problem of selecting the best among $M$ available models, where the identity of the best model is unknown and can change from batch to batch. To isolate the learning challenge, we make several simplifications. We assume each model invocation incurs a unit cost ($c_{m_{t}}=1$), rendering the budget constraint trivial if the per-query budget $b\geq 1$. We also assume that the system can always deploy any of the $M$ models under consideration in a given batch ($M_{\max}\geq M$) and that there are no per-model capacity limits ($\alpha_{m}=1$ for all $m$). The core difficulty then lies in continuously learning and adapting to the best-performing model(s) in each batch.
The proof strategy is to construct a class of adversarial problem instances. We will demonstrate that any online algorithm $\mathcal{A}$ must incur significant regret on at least one instance within this class. This argument adapts a batch-based structure common in analyses of learning problems with non-stationary environments.
Step 1: Construction of Hard Problem Instances.
Let $\epsilon>0$ be a small parameter, which will be determined later. The total time horizon of $T$ queries is partitioned into $N_{B}$ contiguous batches, denoted $\mathcal{B}_{1},\dots,\mathcal{B}_{N_{B}}$. Each batch $j$ consists of $\Delta=T/N_{B}$ queries. For simplicity, we assume $T$ is an integer multiple of $N_{B}$.
For each batch $j\in\{1,\dots,N_{B}\}$, we consider a set of $M$ models available to the algorithm $\mathcal{A}$. These models are indexed $i\in\{1,\dots,M\}$ specifically for the current batch $j$. The performance characteristics of these $M$ models are defined as follows: one model is designated as the “strong” model, and the remaining $M-1$ models are “weak”. Let $s_{j}\in\{1,\dots,M\}$ be the index of this strong model in batch $j$, chosen uniformly at random by the adversary and unknown to the algorithm. The reward $r_{t}$ obtained from selecting a model at query $t\in\mathcal{B}_{j}$ is drawn from a Bernoulli distribution:
  • If model $m_{s_{j}}$ (the strong model for batch $j$, with index $s_{j}$) is chosen, $r_{t}\sim\text{Bernoulli}(\mu_{H}^{(j)})$, where the mean reward is $\mu_{H}^{(j)}=\frac{1}{2}+j\epsilon$.
  • If any other model $m_{i}$ (where $i\in\{1,\dots,M\}$ and $i\neq s_{j}$) from the set of $M$ models for batch $j$ is chosen, $r_{t}\sim\text{Bernoulli}(\mu_{L}^{(j)})$, where the mean reward is $\mu_{L}^{(j)}=\frac{1}{2}+(j-1)\epsilon$.
Rewards from different queries are assumed to be independent. The crucial gap in expected reward between the strong model and any weak model in batch $j$ is $\mu_{H}^{(j)}-\mu_{L}^{(j)}=\epsilon$. To ensure that all mean rewards $\mu_{H}^{(j)}$ and $\mu_{L}^{(j)}$ lie comfortably within the interval $[0,1]$, we impose the condition $N_{B}\epsilon\leq\frac{1}{4}$. This ensures $\mu_{H}^{(N_{B})}=\frac{1}{2}+N_{B}\epsilon\leq\frac{1}{2}+\frac{1}{4}=\frac{3}{4}<1$, and the smallest mean, $\mu_{L}^{(1)}=\frac{1}{2}+(1-1)\epsilon=\frac{1}{2}$, is also valid.
A complete problem instance is characterized by a sequence of strong model indices $(s_{1},s_{2},\dots,s_{N_{B}})$. The actual underlying LLMs corresponding to these indices could differ from batch to batch, but in each batch $j$, the algorithm $\mathcal{A}$ faces a choice among $M$ options with the specified reward structure, and $s_{j}$ is unknown.
Step 2: Per-Batch Regret from Identification Difficulty.
Fix a batch $j\in\{1,\dots,N_{B}\}$, and let $s_{j}^{*}$ denote the true index of the strong model, unknown to algorithm $\mathcal{A}$. To identify $m_{s_{j}^{*}}$ (mean $\mu_{H}^{(j)}$) from the $M-1$ weak models (mean $\mu_{L}^{(j)}$, gap $\epsilon$) with a constant probability $p_{succ}<1$ (e.g., $p_{succ}=3/4$), any algorithm requires a query complexity of $\Omega(M\epsilon^{-2})$ [2, 30]. This implies there is a universal constant $c_{S}>0$ such that at least $c_{S}M\epsilon^{-2}$ queries are necessary to achieve success probability $p_{succ}$.
We set the batch length $\Delta=(c_{S}/2)M\epsilon^{-2}$. Since $c_{S}/2<c_{S}$, this choice of $\Delta$ is insufficient for reliable identification with probability $p_{succ}$. Consequently, algorithm $\mathcal{A}$ fails to identify $m_{s_{j}^{*}}$ within $\Delta$ queries with at least a constant probability $p_{fail}\geq 1-p_{succ}$ (e.g., setting $p_{succ}=3/4$ gives $p_{fail}\geq 1/4$). Let $\mathcal{E}_{fail}$ denote this event of identification failure.
Let $N_{j,weak}$ be the number of queries to weak models in batch $j$. Conditional on $\mathcal{E}_{fail}$, the algorithm lacks knowledge of $s_{j}^{*}$. In such a state of confusion (especially when $M\geq 2$), it is expected to select weak models a significant fraction of the time. For instance, if choices upon failure are made nearly uniformly among the $M$ models, weak models are selected $\frac{M-1}{M}\Delta$ times in expectation. Since $M\geq 2$, $(M-1)/M\geq 1/2$. Thus, we can state that $\mathbb{E}[N_{j,weak}|\mathcal{E}_{fail}]\geq c_{F}\Delta$ for a chosen constant $c_{F}\geq 1/2$. By the law of total expectation, and since $\mathbb{E}[N_{j,weak}|\mathcal{E}_{fail}^{c}]\geq 0$:
$\mathbb{E}[N_{j,weak}]=\mathbb{P}(\mathcal{E}_{fail})\mathbb{E}[N_{j,weak}|\mathcal{E}_{fail}]+\mathbb{P}(\mathcal{E}_{fail}^{c})\mathbb{E}[N_{j,weak}|\mathcal{E}_{fail}^{c}]\geq p_{fail}\cdot c_{F}\Delta.$
Using $p_{fail}\geq 1/4$ and $c_{F}\geq 1/2$, this gives $\mathbb{E}[N_{j,weak}]\geq\frac{1}{4}\cdot\frac{1}{2}\Delta=\frac{\Delta}{8}$. The expected regret in batch $j$ (for a fixed $s_{j}^{*}$, averaged over $\mathcal{A}$’s randomness) is $R_{j}(s_{j}^{*})=\epsilon\mathbb{E}[N_{j,weak}]\geq\epsilon(\Delta/8)$. Since this lower bound does not depend on the specific index $s_{j}^{*}$, and the adversary chooses $s_{j}^{*}$ uniformly, the expected regret in batch $j$ (averaged over $s_{j}^{*}$ and $\mathcal{A}$) is $\mathbb{E}_{\tilde{s}_{j}}[R_{j}]\geq\frac{1}{8}\epsilon\Delta$.
Step 3: Regret Along the Horizon and Parameter Choice.
The expected regret in each batch $j$ is $\mathbb{E}_{\tilde{s}_{j}}[R_{j}]\geq\frac{1}{8}\epsilon\Delta$. The total expected regret over $N_{B}$ batches for algorithm $\mathcal{A}$ is $R_{T}(\mathcal{A})=\sum_{j=1}^{N_{B}}\mathbb{E}_{\tilde{s}_{j}}[R_{j}]\geq N_{B}\frac{1}{8}\epsilon\Delta$. Since $N_{B}=T/\Delta$, we have $R_{T}(\mathcal{A})\geq(T/\Delta)\frac{1}{8}\epsilon\Delta=\frac{1}{8}T\epsilon$.
We have set the batch length $\Delta=(c_{S}/2)M\epsilon^{-2}$ in Step 2. The other primary constraint, from Step 1, is $N_{B}\epsilon\leq 1/4$, which implies $(T/\Delta)\epsilon\leq 1/4$. Substituting $\Delta=(c_{S}/2)M\epsilon^{-2}$ into this constraint:
$\frac{T}{(c_{S}/2)M\epsilon^{-2}}\epsilon\leq\frac{1}{4}\implies\frac{2T\epsilon^{3}}{c_{S}M}\leq\frac{1}{4}\implies\epsilon^{3}\leq\frac{c_{S}M}{8T}.$
To maximize the lower bound on regret $\frac{1}{8}T\epsilon$, we choose $\epsilon$ to be as large as possible, subject to this constraint. We set $\epsilon=\left(\frac{c_{S}M}{8T}\right)^{1/3}$. This choice ensures $\epsilon^{3}=\frac{c_{S}M}{8T}$.
Substituting this $\epsilon$ into the total regret expression:
$R_{T}(\mathcal{A})\geq\frac{1}{8}T\left(\frac{c_{S}M}{8T}\right)^{1/3}=\frac{1}{8}T\frac{c_{S}^{1/3}M^{1/3}}{8^{1/3}T^{1/3}}=\frac{1}{8}\frac{c_{S}^{1/3}}{2}M^{1/3}T^{2/3}=\frac{c_{S}^{1/3}}{16}M^{1/3}T^{2/3}.$
Since $c_{S}>0$ is a universal constant from the identification complexity, let $C=c_{S}^{1/3}/16$. Then $C>0$. Thus, $R_{T}(\mathcal{A})\geq CM^{1/3}T^{2/3}$.
Finally, we verify the conditions on our parameter choices:
  1. 1.
    Value of $\Delta$: $\epsilon^{2}=\left(\frac{c_{S}M}{8T}\right)^{2/3}$. $\Delta=(c_{S}/2)M\epsilon^{-2}=(c_{S}/2)M\left(\frac{c_{S}M}{8T}\right)^{-2/3}=(c_{S}/2)M\frac{(8T)^{2/3}}{(c_{S}M)^{2/3}}$ $=(c_{S}/2)M\frac{8^{2/3}T^{2/3}}{c_{S}^{2/3}M^{2/3}}=(c_{S}/2)M\frac{4T^{2/3}}{c_{S}^{2/3}M^{2/3}}=\frac{c_{S}}{2}\frac{4M^{1/3}T^{2/3}}{c_{S}^{2/3}}=2c_{S}^{1/3}M^{1/3}T^{2/3}$. Let $c_{\Delta}=2c_{S}^{1/3}$. So $\Delta=c_{\Delta}M^{1/3}T^{2/3}$. We need $1\leq\Delta\leq T$ for $\Delta$ to be a valid batch length. $\Delta\leq T\implies c_{\Delta}M^{1/3}T^{2/3}\leq T\implies M^{1/3}\leq T^{1/3}/c_{\Delta}\implies T\geq(c_{\Delta}M^{1/3})^{3}=c_{\Delta}^{3}M$. This implies $T$ must be sufficiently large relative to $M$. $\Delta\geq 1$ generally holds for large $T$ if $M\geq 1$.
  2. 2.
    Constraint $N_{B}\epsilon\leq 1/4$: This constraint was used to determine the choice of $\epsilon$. With $\epsilon^{3}=\frac{c_{S}M}{8T}$:
    $\frac{T\epsilon^{3}}{(c_{S}/2)M}=\frac{T\left(\frac{c_{S}M}{8T}\right)}{(c_{S}/2)M}=\frac{c_{S}M/8}{c_{S}M/2}=\frac{1/8}{1/2}=\frac{1}{4}.$
    Thus, $(T/\Delta)\epsilon=N_{B}\epsilon=1/4$ is satisfied by this construction.
All conditions are met for suitable choices of $T$ relative to $M$, given the universal constant $c_{S}>0$ and our choices for $p_{fail}$ (e.g., $1/4$) and $c_{F}$ (e.g., $1/2$) which determine the factor $1/8$ in the per-batch regret. Thus, for any online routing algorithm $\mathcal{A}$, there exists a problem instance in our constructed class for which its expected regret is bounded below by $R_{T}(\mathcal{A})\geq CM^{1/3}T^{2/3}$ for some constant $C>0$ (specifically $C=c_{S}^{1/3}/16$). For a fixed number of models $M\geq 2$. In this case, the lower bound becomes $\Omega(T^{2/3})$, matching the statement in Theorem 2.
This completes the proof.